### v1

# Norm-equivariant Continuous Metrics Are Adelic

Abstract
Yuan and Zhang showed in [YZ] that any adelic line bundle will induce a norm-equivariant continuous metric of the line bundle on the Berkovich space. We will show the converse that any norm-equivariant continuous metric can be obtained this way. As an application, we will introduce the Monge-Ampère measure on the Berkovich spaces over $\mathbb{Z}$.
1 Introduction
Recently, Yuan and Zhang introduced adelic divisors on quasi-projective varieties in [YZ]. Let $k=\mathbb{Z}$ or a field, and let $U$ be an essentially quasi-projective flat integral scheme over $k$. Let $\widehat{\mathrm{Div}}(U/k)$ be the group of adelic divisors on $U$ defined in [YZ, §2]. It is a reasonable limit of a sequence of divisors on projective models. Roughly speaking, an adelic divisor $\bar{D}\in\widehat{\mathrm{Div}}(U/k)$ is represented by a Cauchy sequence $\bar{D}_{i}\in\widehat{\mathrm{Div}}(X_{i})_{\mathbb{Q}}$ on different projective models $X_{i}$ of $U$, such that all restrictions $\bar{D}_{i}|_{U}$ are equal and it satisfies a Cauchy condition on the boundary. See §3 for more precise definitions.
In [YZ, §3], they also use Berkovich space to define an analytification map of adelic divisors, and then studied the singularity of Green functions along the boundary in the analytic situation. Let $k=\mathbb{Z}$ or a field as before. If $k=\mathbb{Z}$, let $|\cdot|_{\mathrm{Ban}}$ be the usual absolute norm on $\mathbb{Z}$. If $k$ is a field, let $|\cdot|_{\mathrm{Ban}}$ be the trivial norm on $k$. Let $U$ be an essentially quasi-projective flat integral scheme over $k$. Then we have a Berkovich analytic space $U^{\mathrm{an}}=(U/(k,|\cdot|_{\mathrm{Ban}}))^{\mathrm{an}}$ as in [Ber09]. The analytic arithmetic divisors is defined in [YZ, §3].
An analytic arithmetic divisor $\bar{D}=(D,g_{D})$ consists a usual divisor $D\in\mathrm{Div}(U)$ and a continuous function
$g_{D}:U^{\mathrm{an}}\backslash|D|^{\mathrm{an}}\to\mathbb{R},$
such that if $D$ is locally defined by a function $f\in K(U)^{*}$, then $g_{D}(x)+\log|f(x)|$ extends continuously to $|D|^{\mathrm{an}}$ locally. The divisor $\bar{D}=(D,g_{D})$ is called norm-equivariant if for any $x,y\in U^{\mathrm{an}}\backslash|D|^{\mathrm{an}}$ satisfying $|\cdot|_{x}=|\cdot|_{y}^{t}$ for some $t\in(0,\infty)$, then $g_{D}(x)=tg_{D}(y)$.
Let $\widehat{\mathrm{Div}}(U^{\mathrm{an}})_{\mathrm{eqv}}$ be the group of norm-equivariant analytic arithmetic divisors on $U^{\mathrm{an}}$.
In [YZ, §3.3] they define a canonical injective map
$\widehat{\mathrm{Div}}(U/k)\to\widehat{\mathrm{Div}}(U^{\mathrm{an}})_{\mathrm{eqv}}.$
See §3 for details. In particular, let $\widehat{\mathrm{Div}}(U/k)_{b}$ be the group of adelic divisors with integral part zero, then the above map induces an injection
$\widehat{\mathrm{Div}}(U/k)_{b}\to C(U^{\mathrm{an}})_{\mathrm{eqv}}.$
In our paper, we will show that the above analytification maps are surjective when $U$ is quasi-projective.
Theorem 1.1.
(Theorem 3.7) Let $k$ be either $\mathbb{Z}$ or a field. Let $U$ be a quasi-projective flat integral scheme over $k$. Then the above canonical maps
$\widehat{\mathrm{Div}}(U/k)\to\widehat{\mathrm{Div}}(U^{\mathrm{an}})_{\mathrm{eqv}},$
$\widehat{\mathrm{Div}}(U/k)_{b}\to C(U^{\mathrm{an}})_{\mathrm{eqv}}$
are bijective.
We can also get similar results for local fields $K$, but in this case our analytification map is slightly different from Yuan–Zhang’s one.
Let $K$ be either a complete non-archimedean field with a non-trivial discrete valuation, or $K=\mathbb{R}$ or $\mathbb{C}$ with the standard absolute norm. The norm on $K$ is denoted by ||$\cdot$|| ⋅ |.
Let $|\cdot|_{0}$ be the trivial norm on $K$. Define a new norm
$|\cdot|_{\mathrm{Ban}}=\max\{|\cdot|,|\cdot|_{0}\}.$
Then the Berkovich space of the Banach field $(K,|\cdot|_{\mathrm{Ban}})$ is homeomorphic to $[0,1]$, see [Ber09, §2].
Let $U$ be a quasi-projective variety over $K$. If $K$ is non-archimedean, let $\mathscr{O}_{K}$ be the valuation ring of $K$; if $K$ is archimedean, let $\mathscr{O}_{K}=K$. Then Yuan–Zhang defined the group of local adelic divisors $\widehat{\mathrm{Div}}(U/\mathscr{O}_{K})$. Let $U^{\mathrm{an}}=(U/(K,|\cdot|_{\mathrm{Ban}}))^{\mathrm{an}}$ be the Berkovich space of $U$ over $(K,|\cdot|_{\mathrm{Ban}})$, which is also studied in [Ber09, §2]. The group $\widehat{\mathrm{Div}}(U^{\mathrm{an}})_{\mathrm{eqv}}$ is defined similarly.
Theorem 1.2.
The analytification map
$\widehat{\mathrm{Div}}(U/\mathscr{O}_{K})\to\widehat{\mathrm{Div}}(U^{\mathrm{an}})_{\mathrm{eqv}}$
is bijective.
The proof of two theorems are similar.
In §2.2 we discuss the structure of Berkovich spaces. For a quasi-projective variety $U$ over $k$, define the interior part $U^{\beth}\subset U^{\mathrm{an}}$ to be the set of points where the reduction map
$\mathrm{red}:U^{\mathrm{an}}\dashrightarrow U$
is well-defined, and define the boundary part
$U^{b}:=U^{\mathrm{an}}\backslash U^{\beth}.$
Then we will show that the normalized boundary part $\widetilde{U}^{b}:=U^{b}/\!\sim_{\mathrm{eqv}}$ is compact Hausdorff in Theorem 2.7.
This space generalizes the usual Berkovich analytic space as follows. Let $k$ be a field and $|\cdot|_{\mathrm{Ban}}$ be the trivial norm on $k$. Suppose that there is a smooth projective curve $C/k$ and a projective morphism $\pi:X\to C$, such that $U\cong\pi^{-1}(C\backslash\{p\})$. Let $K_{p}$ be the completion of $K(C)$ at $p$, with a fixed non-trivial discrete valuation $|\cdot|_{p}$, then we have canonical homeomorphisms
$U^{b}$ $\cong(U_{K_{p}}/(K_{p},|\cdot|_{p}))^{\mathrm{an}}\times(0,\infty),$
$\widetilde{U}^{b}$ $\cong(U_{K_{p}}/(K_{p},|\cdot|_{p}))^{\mathrm{an}}.$
In the local case when $K$ is a local field with a fixed valuation ||$\cdot$|| ⋅ |, the space $\widetilde{U}^{b}$ contains $(U/(K,|\cdot|))^{\mathrm{an}}$ as an open subset. In particular, we get a non-archimedean compactification of usual quasi-projective analytic space.
In §3 we will review Yuan–Zhang’s adelic divisors and their analytification. Then in §4 we will show that each model divisor with integral part 00 will induce (non-canonically) a continuous function on $\widetilde{U}^{b}$, and all these functions are dense in $C(\widetilde{U}^{b})$. Thus we get Theorem 1.1 and Theorem 1.2.
In §6, we discuss the intersection theory of adelic line bundles. We will define the global Monge-Ampère measure on $\widetilde{U}^{b}$ for integrable adelic line bundles which generalizes the Monge-Ampère measure in the complex case, and the Chambert-Loir measure in the non-archimedean case. Hopefully these spaces and measures are worth studying.
Acknowledgement
I would like to thank my advisor Xinyi Yuan for his valuable advices and encouragement for this paper. I would also than Ruoyi Guo for his helpful feedback on earlier version of this paper.
2 Berkovich spaces
2.1 Preliminaries on Berkovich spaces
Let $(k,|\cdot|_{\mathrm{Ban}})$ be a commutative Banach ring, and $A$ be a $k$-algebra. The Berkovich space $\mathcal{M}(A)$ is the topological space of multiplicative semi-norms $|\cdot|:A\to\mathbb{R}$ whose restriction to $k$ is bounded by $|\cdot|_{\mathrm{Ban}}$. For a point $x\in\mathcal{M}(A)$, the semi-norm is denoted by $|\cdot|_{x}$, and for $f\in A$, the semi-norm $|f|_{x}$ is also denoted by $|f(x)|$, and thus gives a function $|f|:\mathcal{M}(A)\to\mathbb{R}$. The topology on $\mathcal{M}(A)$ is defined as the weakest topology such that for each $f\in A$, the function $|f|$ is continuous.
Note that for each $x\in\mathcal{M}(A)$, the kernel of the semi-norm $|\cdot|_{x}:A\to\mathbb{R}$ is a prime ideal. This induces a continuous contraction map
$\rho:\mathcal{M}(A)\to\operatorname{Spec}A$
sending $x$ to $\rho(x)=\ker|\cdot|_{x}$. Then $|\cdot|_{x}$ can be viewed as a norm on the field $\kappa(\rho(x))$. Let $H_{x}$ be the completion of the field $(\kappa(\rho(x)),|\cdot|_{x})$, called the residue field at $x$. If $H_{x}$ is non-archimedean, let $R_{x}$ be the valuation ring of $H_{x}$.
If $f:A\to B$ is a $k$-algebra homomorphism, $x\in\mathcal{M}(B)$ corresponding to the semi-norm $|\cdot|_{x}$, then $|\cdot|_{x}\circ f:A\to\mathbb{R}$ is a semi-norm on $A$, thus induces a continuous map $f^{*}:\mathcal{M}(B)\to\mathcal{M}(A)$.
In general, if $(k,|\cdot|_{\mathrm{Ban}})$ is a commutative Banach ring, and $X$ is a scheme over $k$, then $X$ is covered by open affine subschemes $\{\operatorname{Spec}A_{i}\}$, and we may glue $\mathcal{M}(A_{i})$ together to get the analytification $X^{\mathrm{an}}$, which is more rigorously written as $(X/(k,|\cdot|_{\mathrm{Ban}}))^{\mathrm{an}}$ if we need to emphasize the norm $|\cdot|_{\mathrm{Ban}}$. If $f:X\to Y$ is a morphism of schemes, then $f$ induces a continuous morphism $f^{\mathrm{an}}:X^{\mathrm{an}}\to Y^{\mathrm{an}}$, thus the analytification map is a functor from the category of schemes over $k$ to the category of topological spaces. We see that the contraction maps $\mathcal{M}(A_{i})\to\operatorname{Spec}A_{i}$ glue to a continuous map $\rho:X^{\mathrm{an}}\to X$.
We also have an inclusion map $i:X\to X^{\mathrm{an}}$ mapping a point $x\in X$ to $(\kappa(x),|\cdot|_{0})$, where $|\cdot|_{0}$ is the trivial norm on $\kappa(x)$. This map is not continuous in general.
In our paper, we only care about the topological space, so we ignore the notion of analytic functions.
By [Ber09, Lemma 1.1, Lemma 1.2], we have:
Proposition 2.1.
(1) If $X$ is of finite type over $k$, then $X^{\mathrm{an}}$ is locally compact.
(2) If $X$ is separated and of finite type over $k$, then $X^{\mathrm{an}}$ is Hausdorff.
(3) If $X$ is projective over $k$, then $X^{\mathrm{an}}$ is compact.
Now assume that $k=\mathbb{Z}$ or a field, and $|\cdot|_{0}$ is the trivial norm on $k$. Let $X$ be a proper scheme over $k$. Let $X^{\mathrm{an}}=(X/(k,|\cdot|_{0}))^{\mathrm{an}}$ be the analytification of $X$. There is a reduction map
$r:X^{\mathrm{an}}\to X$
defined as follows. Since $X$ is proper, by the valuative criterion, for each $x\in U^{\mathrm{an}}$, the natural map $\operatorname{Spec}H_{x}\to X$ extends uniquely to $\operatorname{Spec}R_{x}\to X$. Then $r(x)$ is the image of the closed points of $\operatorname{Spec}R_{x}$ in $X$.
If $k=\mathbb{Z}$ with usual standard absolute norm, then the reduction map is only defined on the subspace of non-archimedean points in $X^{\mathrm{an}}$. This convention is different from Yuan–Zhang’s one in [YZ, §3]. We will deal with it more carefully in §2.2.
2.2 Structure of Berkovich Spaces
In this section we study the more detailed structure of the Berkovich spaces. First we list all possible choices of the Banach ring $(k,|\cdot|_{\mathrm{Ban}})$ used in our paper. For any ring $k$, the semi-norm $|\cdot|_{0}$ will denote the trivial semi-norm on $k$.
Assumption 2.2.
In this section, a Banach ring $(k,|\cdot|_{\mathrm{Ban}})$ will be one of the following.
  1. (1)
    Non-archimedean global case: $k$ is any noetherian integral domain with the trivial norm $|\cdot|_{\mathrm{Ban}}=|\cdot|_{0}$.
  2. (2)
    Non-archimedean local case: $k$ is a field with a complete discrete non-trivial valuation ||$\cdot$|| ⋅ |. Then $|\cdot|_{\mathrm{Ban}}=\max\{|\cdot|,|\cdot|_{0}\}$.
  3. (3)
    Archimedean local case: $k=\mathbb{R}$ or $\mathbb{C}$, and let ||$\cdot$|| ⋅ | be the standard absolute value. Then $|\cdot|_{\mathrm{Ban}}=\max\{|\cdot|,|\cdot|_{0}\}$.
  4. (4)
    Global case: $k=\mathbb{Z}$, and the norm $|\cdot|_{\mathrm{Ban}}$ is the standard absolute norm on $\mathbb{Z}$.
Note that in all cases, the trivial semi-norm $|\cdot|_{0}$ is bounded by $|\cdot|_{\mathrm{Ban}}$. The first two cases are the non-archimedean cases, and the last two contain both archimedean and non-archimedean cases.
Let $(k,|\cdot|_{\mathrm{Ban}})$ be as in Assumption (2.2). Let $U$ be a quasi-projective flat variety over $k$, and let $U^{\mathrm{an}}=(U/(k,|\cdot|_{\mathrm{Ban}}))^{\mathrm{an}}$. For two points $x,y\in U^{\mathrm{an}}$, we say that $x,y$ are norm-equivalent, denoted by $x\sim_{\mathrm{eqv}}y$, if $\rho(x)=\rho(y)$ and $|\cdot|_{y}=|\cdot|_{x}^{t}$ for some positive real number $t$. This is an equivalence relation, and we are interested in $X^{\mathrm{an}}/\!\sim_{\mathrm{eqv}}$. In general, it is almost never a Hausdorff space, thus we use a subspace of $U^{\mathrm{an}}$ instead. We first define the reduction map carefully.
Definition 2.3.
The interior part of $U^{\mathrm{an}}$, denoted by $U^{\beth}\subset U^{\mathrm{an}}$, is the set of all points in $U^{\mathrm{an}}$ such that $|a|_{x}\leq 1$ for all $a\in k$(in particular, $x$ is a non-archimedean point, and there is a morphism $\operatorname{Spec}R_{x}\to\operatorname{Spec}k$), and that there exists a morphism $\varphi:\operatorname{Spec}R_{x}\to X$ such that the commutative diagram
commutes.
The boundary part $U^{b}$ is defined as $U^{b}=U^{\mathrm{an}}\backslash U^{\beth}$, and the normalized boundary part is $\widetilde{U}^{b}/\sim_{\mathrm{eqv}}$.
Since $U$ is separated over $k$, there is at most one extension $\varphi$. The reduction map $r:U^{\beth}\to U$ is defined similarly as the projective case. For $x\in U^{\beth}$, $r(x)$ is the image of the unique closed points of $\operatorname{Spec}R_{x}$ under the morphism $\varphi$.
Proposition 2.4.
Let $k$ be as in Assumption 2.2. Then the interior part $U^{\beth}$ is compact Hausdorff.
Proof.
Let $X$ be a projective model of $U$, and $X^{\mathrm{an}}=(X/(k,|\cdot|_{\mathrm{Ban}}))^{\mathrm{an}}$. Since $X$ is proper over $k$, $X^{\beth}$ is just the set of points in $X^{\mathrm{an}}$ such that $|a|_{x}\leqslant 1$ for all $a\in k$. Equivalently, we have
$X^{\beth}\cong(X/(k,|\cdot|_{0}))^{\mathrm{an}},$
and hence $X^{\beth}$ is compact Hausdorff.
By [Ber90], the reduction map $r:X^{\beth}\to X$ is anti-continuous, i.e., the inverse image of an open subset is closed. Thus $U^{\beth}=r^{-1}(U)$ is a closed subspace of $X^{\beth}$, hence it is compact Hausdorff. ∎
Remark 2.5.
Here we explain the notations. If $k$ is a field with trivial norm, in [Thu, §1], Thuillier defined an analytic space $\mathfrak{X}^{\beth}$ for any locally algebraic formal $k$-scheme $\mathfrak{X}$. Then ${U}^{\beth}$ is the analytification of the formal scheme ${U}$(which is just the scheme $U$ with the discrete topology on the structure sheaf).
Let $X$ be a projective model of $U$, and let $\mathfrak{X}$ be the formal completion of $X$ along $D=X\backslash U$. Then $\mathfrak{X}^{\beth}=U^{b}\cup D^{\mathrm{an}}$, and $U^{b}=\mathfrak{X}_{\eta}$ is equal to the generic fiber of the formal scheme $\mathfrak{X}$ in [Thu]. We choose this notation just because it only depends on $U$, not on the projective model, and we want to emphasize that $U^{\mathrm{an}}=U^{\beth}\cup U^{b}$.
First we give a local structure of $\widetilde{U}^{b}$ when $k$ is a field with trivial norm.
Proposition 2.6.
If $k$ is a field with trivial norm, let $k((t))$ be the complete discrete valuation field with a fixed norm ||$\cdot$|| ⋅ |. Then $\widetilde{U}^{b}$ is locally homeomorphic to analytification of some scheme of finite type over $k((t))$.
Proof.
Let $X$ be any projective model of $U$ such that $D=X\backslash U$ is a divisor on $X$.
Take any $p\in X\backslash U$, and any open affine neighborhood $\operatorname{Spec}A$ of $p$. Suppose that $D$ is locally defined by $f\in A$ on $\operatorname{Spec}A$. Then
[…]
and $U^{b}$ is covered by such open subspaces.
If we fix $r\in(0,1)$, then the quotient space $\left(U^{b}\cap(\operatorname{Spec}A)^{\mathrm{an}}\right)/\!\sim_{\mathrm{eqv}}$ is homeomorphic to the space
$\{x\in\mathcal{M}(A):|a|_{x}\leqslant 1,\forall a\in A;\ |f|_{x}=r\},$
which is homeomorphic to the $k((t))$-analytic space $\operatorname{Spec}A\widehat{\otimes}_{k[t]}k((t))$ where the morphism $k[t]\to A$ is given by $t\mapsto f$. ∎
The above homeomorphic depends on the choice of $f$, and we cannot expect to have a global one, let along a canonical one. If there is a projective morphism $\pi:X\to C$ where $C$ is a smooth projective curve, and $D=\pi^{-1}p$ for some closed point $p\in C$, then $\widetilde{U}^{b}$ is homeomorphic to the $\widehat{K(C)_{p}}$-analytic space $(X\times_{C}\operatorname{Spec}\widehat{K(C)}_{p})^{\mathrm{an}}$.
Theorem 2.7.
Let $k$ be as in Assumption 2.2, and $U$ be a quasi-projective integral flat scheme over $k$. Then the normalized boundary part $\widetilde{U}^{b}$ is compact Hausdorff.
We will prove Theorem 2.7 case by case.
Non-archimedean global case
In this subsection we prove Theorem 2.7 in the non-archimedean global case when $k$ is any noetherian integral domain with the trivial norm $|\cdot|_{\mathrm{Ban}}=|\cdot|_{0}$.
Let $U$ be a quasi-projective integral flat scheme over $k$. Fix a projective model $X/k$ of $U$, and let $D=X\backslash U$, with the reduced subscheme structure. Then the reduction map $r:X^{\mathrm{an}}\to X$ is defined on the whole space $X^{\mathrm{an}}$.
Proof of Theorem 2.7.
First we show the Hausdorffness of $\widetilde{U}^{b}=U^{b}/\!\sim_{\mathrm{eqv}}$.
Let $x,y\in U^{b}$ be two points which are not norm-equivalent. Then $r(x),r(y)$ are two points in $D$(possibly equal). Since $X$ is projective, there exists an open affine subscheme $\operatorname{Spec}A\subset X$ containing $r(x)$ and $r(y)$. Then $\operatorname{Spec}A$ also contains $\rho(x)$ and $\rho(y)$, and $x,y$ give two norm $|\cdot|_{x},|\cdot|_{y}$ on $A$.
Since $r(x),r(y)\in\operatorname{Spec}A$, we have $|a|_{x}\leqslant 1$, $|a|_{y}\leqslant 1$ for all $a\in A$.
Note that $\rho(y)\notin\overline{\{r(x)\}}$, as $\rho(y)\notin D$ and $r(x)\in D$. Thus we have
$\{a\in A:|a|_{x}<1\}\not\subset\{a\in A:|a|_{y}=0\}.$
By prime avoidance, two prime ideals $\{a\in A:|a|_{x}=0\}$ and $\{a\in A:|a|_{y}=0\}$ cannot cover the ideal $\{a\in A:|a|_{x}<1\}$. Thus there exists $f_{x}\in A$ such that $10<|f_{x}|_{x}<$ 1 and $|f_{x}|_{y}\neq 0$.
Similarly, take $f_{y}\in A$ such that $10<|f_{y}|_{y}<$ 1 and $|f_{y}|_{x}\neq 0$. Let $f=f_{x}f_{y}$. Then we have $10<|f|_{x}<$ 1 and $10<|f|_{y}<$ 1.
Since $x,y$ are not norm-equivalent, switch $x,y$ if necessary, we may find $g\in A$, such that
$\frac{\log|g(x)|}{\log|f(x)|}<\frac{\log|g(y)|}{\log|f(y)|}.$
Take any $a\in\mathbb{R}$ such that
$\frac{\log|g(x)|}{\log|f(x)|}<a<\frac{\log|g(y)|}{\log|f(y)|},$
Define two open subset of $(\operatorname{Spec}A)^{\mathrm{an}}$ as
$V_{1}=\{v\in(\operatorname{Spec}A)^{\mathrm{an}}:f(v)\neq 0,1;\ \frac{\log|g(v)|}{\log|f(v)|}<a\}$
$V_{2}=\{v\in(\operatorname{Spec}A)^{\mathrm{an}}:f(v)\neq 0,1;\ \frac{\log|g(v)|}{\log|f(v)|}>a\}$
Here we assume that $\log 0=-\infty<a$ for all real number $a$. Then $V_{1}$, $V_{2}$ are two disjoint open subsets of $(\operatorname{Spec}A)^{\mathrm{an}}$ which are stable under norm-equivalence relation and separate $x,y$. Thus $\widetilde{U}^{b}$ is Hausdorff.
For the compactness of $\widetilde{U}^{b}$, since $X^{\mathrm{an}}$ is compact Hausdorff, it is a $T_{4}$-topological space. As $U^{\beth}$, $D^{\mathrm{an}}$ are both closed in $X^{\mathrm{an}}$, there exist disjoint open subsets $W_{1}\supset U^{\beth}$ and $W_{2}\supset D^{\mathrm{an}}$. Let $W_{3}=X^{\mathrm{an}}\backslash(W_{1}\cup W_{2})$, then $W_{3}$ is closed in $X^{\mathrm{an}}$, and hence compact.
We show that $W_{3}\to U^{b}\to\widetilde{U}^{b}$ is surjective. Given any $x\in U^{b}$, choose an affine open subset $\operatorname{Spec}A$ containing $r(x)$, then $x$ corresponding to a norm $|\cdot|_{x}$ on $A$. Define the line $L:[0,\infty]\to X^{\mathrm{an}}$ by
$L(t)=\begin{cases}i(\rho(x)),&\text{if }t=0;\\ |\cdot|_{x}^{t},&\text{if }t\in(0,\infty);\\ i(r(x)),&\text{if }t=\infty.\end{cases}$
Then we can check that $L$ is continuous. Since $L(0)\in W_{1}$ ,and $L(\infty)\in W_{2}$, and $[0,\infty]$ is connected, $L^{-1}(W_{1})$ and $L^{-1}(W_{2})$ cannot cover $[0,\infty]$. Thus there exist $y\in W_{3}$, and $y$ is norm-equivalent to $x$. Thus $W_{3}\to\widetilde{U}^{b}$ is surjective.
Now $\widetilde{U}^{b}$ is a continuous image of a compact space, and hence it is compact. ∎
Non-archimedean local case
If $K$ is a non-archimedean field with a non-trivial discrete valuation, any quasi-projective integral scheme over $K$ is also quasi-projective over $\mathscr{O}_{K}$. Now let $U$ be a quasi-projective integral scheme over $K$. We have two analytic spaces
$U_{\mathrm{Ban}}^{\mathrm{an}}=(U/(K,|\cdot|_{\mathrm{Ban}}))^{\mathrm{an}},\qquad U_{\mathscr{O}_{K}}^{\mathrm{an}}=(U/(\mathscr{O}_{K},|\cdot|_{0}))^{\mathrm{an}}.$
Note that $U_{1}^{\mathrm{an}}$ is in fact a closed subspace of $U_{2}^{\mathrm{an}}$. Indeed, if $\pi$ is a uniformizer of $K$, then we have a homeomorphism $\mathcal{M}(\mathscr{O}_{K},|\cdot|_{0})\cong[0,\infty]$ given by $|\cdot|_{x}\mapsto-\log|\pi|_{x}$. Then we have the structure morphism $f:U_{2}^{\mathrm{an}}\to\mathcal{M}(\mathscr{O}_{K},|\cdot|_{0})\cong[0,\infty]$, and $U_{\mathrm{Ban}}^{\mathrm{an}}\cong f^{-1}([0,|\pi|])$. Thus we may identify $U_{\mathrm{Ban}}^{\mathrm{an}}$ with a closed subspace of $U_{2}^{\mathrm{an}}$, and we have a canonical homeomorphism $U_{\mathrm{Ban}}^{\beth}\cong U_{2}^{\beth}$ under this identification.
Since the image of $f$ is contained in $[0,\infty)$, any points in $U_{2}^{\mathrm{an}}$ is norm-equivalent to some point in $U_{1}^{\mathrm{an}}$. Thus the spaces $\widetilde{U}^{b}$ computed in both spaces are canonically homeomorphic, and hence it is compact Hausdorff.
Archimedean local case and the global case
Now let $k=\mathbb{C}$ be the field of complex numbers, and ||$\cdot$|| ⋅ | be the usual absolute norm on $\mathbb{C}$.
Now let $U$ be a quasi-projective variety over $\mathbb{C}$, and $X$ be a fixed projective model of $U$. Let $U^{\mathrm{an}}=(U/(\mathbb{C},|\cdot|_{\mathrm{Ban}}))^{\mathrm{an}}$ be the Berkovich space with structure morphism $\pi:U^{\mathrm{an}}\to\mathcal{M}(\mathbb{C},|\cdot|_{\mathrm{Ban}})\cong[0,1]$ and let $U_{t}^{\mathrm{an}}=\pi^{-1}(t)$ be the fiber over $t$. Similarly, define $X^{\mathrm{an}}$ and $X^{\mathrm{an}}_{t}$ for $t\in[0,1]$.
Points in $U_{0}^{\mathrm{an}}$ are non-archimedean points, and all others are archimedean. The fiber $U_{1}^{\mathrm{an}}$ are just usual complex analytification of $U$.
We first show that $\pi^{-1}((0,1])$ is homeomorphic to $U_{1}^{\mathrm{an}}\times(0,1]$. This is proved in [Ber09, Lemma 2.1]. He implicitly used the following lemma.
Lemma 2.8.
[…]
Proof.
We only need to check the triangle inequality. By direct computation, we get:
$\|x+y\|^{n}$ $\leqslant\sum_{i=0}^{n}\left|{n\choose i}\right|^{t}\|x\|^{i}\|y\|^{n-i}$
$\leqslant(n+1)\left(\sum_{i=0}^{n}\left|{n\choose i}\right|\|x\|^{\frac{i}{t}}\|y\|^{\frac{n-i}{t}}\right)^{t}$
$=(n+1)\left(\|x\|^{\frac{1}{t}}+\|y\|^{\frac{1}{t}}\right)^{nt}$
Take $nt$-th roots, and let $n\to\infty$, we get
$\|x+y\|^{\frac{1}{t}}\leqslant\|x\|^{\frac{1}{t}}+\|y\|^{\frac{1}{t}}$
In particular, we see that $U_{1}^{\mathrm{an}}$ is an open subset of $\widetilde{U}^{b}$.
Now it’s ready to show that $\widetilde{U}^{b}$ is compact Hausdorff.
Proof of Theorem 2.7.
We first show Hausdorffness.
Let $x,y$ be two points in $U^{b}$ which are not norm-equivalent. If $x,y$ are both archimedean points, it follows that $U_{1}^{\mathrm{an}}$ is open in $U^{b}/\!\sim_{\mathrm{eqv}}$, and $U_{1}^{\mathrm{an}}$ is Hausdorff.
If $x$ is archimedean and $y$ is non-archimedean, we choose an open affine subset $\operatorname{Spec}A\subset X$ containing $\rho(x)$ and $r(y)$. Then we may find $f\in A$ such that $10<|f|_{x}<$ 1 and $|f|_{y}\neq 0$ by prime avoidance. By multiplying a nonzero $z\in\mathbb{C}$, we may assume that $|f|_{x}\neq 0,1$ and $|f|_{y}\neq 0,1$ hold simultaneously.
If $x,y$ are both non-archimedean, similarly as the non-archimedean global case, we can also find $f$ such that $|f|_{x}\neq 0,1$ and $|f|_{y}\neq 0,1$ hold simultaneously. The rest part of the proof are similar as the non-archimedean case.
To show the compactness, choose a projective model $X$ and let $D=X\backslash U$. Let $Y=D^{\mathrm{an}}\cup X_{1}^{\mathrm{an}}$, which is closed subset in $X^{\mathrm{an}}$. Then $U^{\beth},Y$ are both compact. Take disjoint open neighborhood $W_{1}\supset U^{\beth}$, $W_{2}\supset D^{\mathrm{an}}$, and let $W_{3}=X^{\mathrm{an}}\backslash(W_{1}\cup W_{2})$. Then similarly $W_{2}$ maps surjectively onto $\widetilde{U}^{b}$, thus $\widetilde{U}^{b}$ is compact. ∎
The case $k=\mathbb{Z}$ with the standard absolute norm can be proved similarly as the archimedean case.
3 Adelic divisors on Berkovich spaces
The adelic divisors and line bundles on quasi-projective varieties were introduced by Yuan–Zhang in [YZ, §2]. We just briefly review their definitions.
3.1 Adelic divisors
Let $k=\mathbb{Z}$ or a field, and let $X$ be a projective flat variety over $k$.
If $k=\mathbb{Z}$, an arithmetic divisor on $X$ is a pair $(D,g_{D})$, where $D$ is a Cartier divisor on $X$, and $g_{D}$ is a continuous real-valued conjugation-invariant function on $(X\backslash D)(\mathbb{C})$(with respect to the analytic topology), such that if $D$ is locally defined by $f$, then $g_{D}(x)+\log|f(x)|$ extends continuously to $D(\mathbb{C})$. Equivalently, $e^{-g_{D}}$ induces a conjugation-invariant metric on the line bundle $\mathscr{O}(D)$ over $X(\mathbb{C})$, such that the canonical section $1\in\mathscr{O}(D)$ has the norm $\|1\|(x)=e^{-g_{D}(x)}$. The arithmetic divisor $(D,g_{D})$ is called effective if $D$ is effective and $g_{D}$ is non-negative.
If $k$ is a field, by abuse of notation, an arithmetic divisor on $X$ is just a usual divisor on $X$. The set of arithmetic divisors on $X$ is denoted by $\widehat{\mathrm{Div}}(X)$.
Now let $U$ be a quasi-projective flat variety over $k$, and let $X$ be a projective model of $U$. Define the group of $(\mathbb{Q},\mathbb{Z})$-divisors as
$\widehat{\mathrm{Div}}(X,U):=\widehat{\mathrm{Div}}(X)_{\mathbb{Q}}\times_{\mathrm{Div}(U)_{\mathbb{Q}}}\mathrm{Div}(U).$
Thus a $(\mathbb{Q},\mathbb{Z})$-divisor $(\bar{D},D^{\prime})$ consists of an arithmetic $\mathbb{Q}$-divisor $\bar{D}\in\widehat{\mathrm{Div}}(X)_{\mathbb{Q}}$ and a usual divisor $D^{\prime}\in\mathrm{Div}(U)$ such that their image in $\widehat{\mathrm{Div}}(U)_{\mathbb{Q}}$ are the same. By abuse of notations, we just write $\bar{D}\in\widehat{\mathrm{Div}}(X,U)$, and call $D^{\prime}$ the integral part of $\bar{D}$.
The group of model divisors on $U$ is defined as
$\widehat{\mathrm{Div}}(U)_{\mathrm{mod}}:=\lim_{\begin{subarray}{c}\longrightarrow\\ X\end{subarray}}\widehat{\mathrm{Div}}(X,U),$
where $X$ ranges over all projective model of $U$, and the morphisms are pull-back of divisors.
There is a boundary norm on $\widehat{\mathrm{Div}}(U)_{\mathrm{mod}}$, which more rigorously is an extended norm, i.e. it allows that $|D|=\infty$ for some $D$. Choose a fixed projective model $X$ such that $Z=X\backslash U$ is a divisor. If $k=\mathbb{Z}$, we further choose a positive Green function $g_{Z}$, and let $\bar{Z}=(Z,g_{Z})$. Let $\bar{D}\in\widehat{\mathrm{Div}}(U)_{\mathrm{mod}}$. The boundary norm $\|\cdot\|_{\bar{Z}}$ is defined as
$\|\bar{D}\|_{\bar{Z}}=\inf\{\varepsilon\in\mathbb{Q}_{>0}:-\varepsilon\bar{Z}\leqslant\bar{D}\leqslant\varepsilon\bar{Z}\},$
where the partial order is given by effectivity of divisors. If the integral part of $\bar{D}$ is non-zero, such an $\varepsilon$ does not exist, and we assume that $\inf\emptyset=\infty$, thus $\|\bar{D}\|_{\bar{\mathbb{Z}}}=\infty$ in this case.
Different boundary divisors give different boundary norms on $\widehat{\mathrm{Div}}(U)_{\mathrm{mod}}$, but they all induce the same boundary topology. The group of adelic divisor $\widehat{\mathrm{Div}}(U/k)$ is defined as the completion of $\widehat{\mathrm{Div}}(U)_{\mathrm{mod}}$ with respect to any boundary norm.
3.2 Analytification of Adelic Divisors
Let $k$ be a commutative Banach ring which is also an integral domain, and $X$ is an integral scheme over $k$. An arithmetic divisor $\bar{D}$ on $X^{\mathrm{an}}$ is a pair $(D,g_{D})$, where $D$ is a usual divisor on $X$, and $g_{D}$ is a continuous function on $X^{\mathrm{an}}\backslash D^{\mathrm{an}}$, such that if $D$ is locally defined by a function $f$, then $g_{D}(x)+\log|f(x)|$ extends continuously to $D^{\mathrm{an}}$.
The arithmetic divisor $\bar{D}$ is called norm-equivariant if for any $x_{1},x_{2}\in X^{\mathrm{an}}\backslash D^{\mathrm{an}}$ satisfying $|\cdot|_{x_{1}}=|\cdot|^{t}_{x_{2}}$ for some $0<t<\infty$, we have that $g(x_{1})=tg(x_{2})$.
Let $\widehat{\mathrm{Div}}(X^{\mathrm{an}})_{\mathrm{eqv}}$ be the group of norm-equivariant arithmetic divisors. Yuan–Zhang defined an analytification map of adelic divisors.
Let $k$ be either $\mathbb{Z}$ or a field. If $k=\mathbb{Z}$, then $|\cdot|_{\mathrm{Ban}}$ is the usual absolute norm on $\mathbb{Z}$. If $k$ is a field, then $|\cdot|_{\mathrm{Ban}}$ is the trivial norm on $k$.
Proposition 3.1.
[YZ, Prop. 3.3.1] Let $X$ be a flat and essentially quasi-projective integral scheme over $k$, and let $X^{\mathrm{an}}=(X/(k,|\cdot|_{\mathrm{Ban}}))^{\mathrm{an}}$. Then there is a canonical injection map
$\widehat{\mathrm{Div}}(X/k)\to\widehat{\mathrm{Div}}(X^{\mathrm{an}})_{\mathrm{eqv}}.$
We briefly discuss their construction. First assume that $X$ is projective. If $k$ is a field, then there is a reduction map $r:X^{\mathrm{an}}\to X$. Let $D$ be a usual divisor on $X$, and let $x\in X^{\mathrm{an}}$. Take any open affine neighborhood $U=\operatorname{Spec}A$ of $r(x)$, such that $D|_{U}$ is locally defined by $f\in k(U)^{*}$. Define $\tilde{g}(x)=-\log|f(x)|$. It is independent of the choice $(U,f)$. Then $(D,\tilde{g})$ is the analytification of $D\in\widehat{\mathrm{Div}}(X/k)$.
If $k=\mathbb{Z}$, let $\bar{D}=(D,g)\in\widehat{\mathrm{Div}}(X)$. The analytification $(D,\tilde{g})$ is defined as follows. The restriction of $\tilde{g}$ to the non-archimedean part of $X^{\mathrm{an}}$ are defined in the same way as above. The archimedean part of $X^{\mathrm{an}}$ is homeomorphic to $X(\mathbb{C})\times(0,1]$ modulo complex conjugaion, thus the restriction of $\tilde{g}$ to the archimedean part is uniquely determined by $g$.
Let $X$ be a fixed projective model of $U$, such that $D=X\backslash U$ is a divisor. Let $\bar{D}$ be a boundary divisor defining the boundary norm, which induces a Green function $g_{\bar{D}}$ on $U^{\mathrm{an}}$. By the above construction, we have the following result.
Lemma 3.2.
For any $x\in U^{\mathrm{an}}$, we have that $g_{\bar{D}}(x)\geqslant 0$, and $g_{\bar{D}}(x)=0$ if and only if $x\in U^{\beth}$.
In general, if $U$ is quasi-projective or essentially quasi-projective, one may take limit on both sides. See [YZ, §3] for details.
Definition 3.3.
Let $X,U,\bar{D}$ as above. Let $\delta_{\bar{D}}(U)$ be the subspace of $U^{\mathrm{an}}$ defined by
$\delta_{\bar{D}}(U)=\{x\in U^{b}:g_{\bar{D}}(x)=1\}\cup\{x\in U^{b}:g_{\bar{D}}(x)=\max_{y\sim_{\mathrm{eqv}}x}{g_{\bar{D}}(y)}\leqslant 1\}.$
Note that the last set is empty unless $U^{\mathrm{an}}$ contains archimedean part. In particular, if $(k,|\cdot|_{\mathrm{Ban}})$ is non-archimedean, then
$\delta_{\bar{D}}(U)=\{x\in U^{\mathrm{an}}:g_{\bar{D}}(x)=1\}.$
Theorem 3.4.
The space $\delta_{\bar{D}}(U)$ is compact Hausdorff.
Proof.
We fix a projective model $X$ with a fixed boundary divisor $\bar{D}=(D,g_{\bar{D}})\in\widehat{\mathrm{Div}}(X)$. If $D$ is locally defined by a function $f$, then $g_{\bar{D}}(x)+\log|f(x)|$ extends continuously to $D^{\mathrm{an}}$. Thus $g_{\bar{D}}:U^{\mathrm{an}}\to\mathbb{R}$ extends continuously to $\bar{g}_{\bar{D}}:X^{\mathrm{an}}\to\mathbb{R}\cup\{\infty\}$.
Now we have equality of two subspaces
$\{x\in U^{b}:g_{\bar{D}}(x)=1\}=\bar{g}_{\bar{D}}^{-1}(1),$
hence it is compact. The subspace $\{x\in U^{b}:g_{\bar{D}}(x)=\displaystyle\max_{y\sim_{\mathrm{eqv}}x}{g_{\bar{D}}(y)}\leqslant 1\}$ is homeomorphic to the space
$\{x\in U/(\mathbb{C},|\cdot|)^{\mathrm{an}}:g_{\bar{D}}(x)\leqslant 1\},$
which is also compact Hausdorff. ∎
Corollary 3.5.
The composition
$\delta_{\bar{D}}(U)\to U^{b}\to\widetilde{U}^{b}$
is a homeomorphism.
Proof.
It is a continuous bijection from a compact space to a Hausdorff space, hence it is a homeomorphism. ∎
Corollary 3.6.
If $(k,|\cdot|_{\mathrm{Ban}})$ is non-archimedean, then we have homeomorphism
$U^{b}\cong\widetilde{U}^{b}\times\mathbb{R}$
The space $\delta_{\bar{D}}(U)$ is easier to study, but it depends on the choice of the boundary divisor $\bar{D}$. The space $\widetilde{U}^{b}$ is intrinsic. However, a Green function $g_{\bar{D}}$ does not induce any reasonable function on $\widetilde{U}^{b}$. We can only consider the function $g_{\bar{D}_{1}}/g_{\bar{D}_{2}}$, so we still need to choose a boundary divisor.
Our main theorem is as follows.
Theorem 3.7.
(Theorem 1.1) Let $k$ be either $\mathbb{Z}$ or a field. Let $U$ be a flat quasi-projective integral scheme over $k$. Then the above canonical map
$\widehat{\mathrm{Div}}(U/k)\to\widehat{\mathrm{Div}}(U^{\mathrm{an}})_{\mathrm{eqv}}$
is an isomorphism.
Let $K$ be a field complete with respect to a non-trivial discrete valuation ||$\cdot$|| ⋅ |. If $K$ is non-archimedean, let $O_{K}$ be the valuation ring of $K$. If $K$ is archimedean, let $O_{K}=K$. Let $U$ be a quasi-projective variety over $K$.
Yuan and Zhang also define the group of adelic divisors $\widehat{\mathrm{Div}}(U/\mathscr{O}_{k})$ in the this local case, and define injections $\widehat{\mathrm{Div}}(U/\mathscr{O}_{K})\to\widehat{\mathrm{Div}}((U/(K,|\cdot|))^{\mathrm{an}})$. However, to get a similar results as in the global case, we need a different analytification.
Let $(K,|\cdot|)$ as above, and let $|\cdot|_{0}$ be the trivial norm on $K$. Define a new norm $|\cdot|_{\mathrm{Ban}}=\max\{|\cdot|,|\cdot|_{0}\}$ on $K$, and let $U^{\mathrm{an}}=(U/(K,|\cdot|_{\mathrm{Ban}}))^{\mathrm{an}}$. Similarly as in the global case, there is a natural analytification map $\widehat{\mathrm{Div}}(U/\mathscr{O}_{K})\to\widehat{\mathrm{Div}}(U^{\mathrm{an}})_{\mathrm{eqv}}$.
Theorem 3.8.
Notations are as above. Then the natural map $\widehat{\mathrm{Div}}(U/\mathscr{O}_{K})\to\widehat{\mathrm{Div}}(U^{\mathrm{an}})_{\mathrm{eqv}}$ is an isomorphism.
4 Proof of the Main Theorems
We first prove the case that $k$ is a field (with the trivial norm), which works equally well for non-archimedean fields. Then we prove the archimedean case and the global case.
Non-archimedean case
First we consider the non-archimedean case, where $k$ is any noetherian integral domain with the trivial norm ||$\cdot$|| ⋅ |.
Let $X$ be a fixed projective model of $U/\mathscr{O}_{k}$, such that $D=X\backslash U$ is a divisor. Then each ideal sheaf $\mathcal{I}\subset\mathscr{O}_{X}$ defines a closed subscheme $Y\subset X$.
Now suppose that $|Y|\subset|D|$. Such ideal sheaves are called boundary ideal sheaves. Then $\mathrm{Bl}_{Y}X$ ia also a projective model of $U$, and the exceptional divisor $E_{Y}$ is a model divisor in $\mathrm{Div}(U)_{\mathrm{mod}}$. The analytification map gives an arithmetic divisor $(0,g_{Y})$ on $U^{\mathrm{an}}$ with integral part zero, and a continuous function $g_{Y}(x)$ on $U^{\mathrm{an}}$.
The function $g_{Y}$ has the form
$g_{Y}(x)=\min\{-\log|a(x)|:a\in\mathcal{I}_{r(x)}\}.$
It is identically zero on $U^{\beth}$, and is always norm-equivariant and non-negative. If further that $Y=D$, we see that $g_{D}(x)$ is strictly positive on $U^{b}$.
Since $g_{Y}(x)$ and $g_{D}(x)$ are both norm-equivariant, the quotient $h_{Y}(x)=\frac{g_{Y}(x)}{g_{D}(x)}$ is a norm-invariant continuous function on $U^{b}$, thus it is the pull-back of a continuous function $h_{Y}$ on $\widetilde{U}^{b}$. Such a function is called a model function on $\widetilde{U}^{b}$ associated to the closed subscheme $Y$. If $Y$ is defined by the boundary ideal sheaf $\mathcal{I}$, we may also write $g_{\mathcal{I}}=g_{Y}$ and $h_{\mathcal{I}}=h_{Y}$.
The constant function $1=h_{D}$ is a model function. Since we have
$h_{\mathcal{I}_{1}\mathcal{I}_{2}}=h_{\mathcal{I}_{1}}+h_{\mathcal{I}_{2}}$
and
$h_{\mathcal{I}_{1}+\mathcal{I}_{2}}=\min\{h_{\mathcal{I}_{1}},h_{\mathcal{I}_{2}}\},$
the space of model functions are closed under addition and taking minimum. To show that the $\mathbb{Q}$-linear combination of model functions associated to the boundary ideal sheaf is dense in the Banach space of continuous functions on $\widetilde{U}^{b}$, we need to show that model functions separate points.
Proposition 4.1.
Model functions on $\widetilde{U}^{b}$ separate points.
Proof.
The proof here is almost the same as in [BFJ, Prop. 2.2]
Let $x,y\in U^{b}$ are two points which are not norm-equivalent. Choose a suitable $x^{\prime}$ which is norm-equivalent to $x$ if necessary, we may assume that $g_{D}(x)=g_{D}(y)$, and still $x\neq y$.
If $r(x)\neq r(y)$, we may assume that $r(y)\notin\overline{r(x)}$. Let $Y=\overline{r(x)}$. Then $h_{Y}(x)>0$ and $h_{Y}(y)=0$, thus $h_{Y}$ separates $x,y$.
Now assume that $r(x)=r(y)$. Take an affine open neighborhood $\operatorname{Spec}A$ of $r(x)$. Since $x\neq y$, there exists $f\in A$ such that $|f(x)|\neq|f(y)|$. Consider the ideal sheaf $\mathscr{O}_{\operatorname{Spec}A}\cdot f$, which can be extended to an ideal sheaf $\mathcal{I}$ on $X$. For each positive integer $m$, the ideal sheaf $\mathcal{I}_{m}=\mathcal{I}+(\mathcal{I}_{D})^{m}$ is a boundary ideal sheaf on $X$, and we have
$h_{\mathcal{I}_{m}}=\min\{\frac{-\log|f|}{g_{D}},m\}$
at $x$ and $y$, hence it separates $x,y$ for $m$ large. ∎
By Boolean ring version of Stone-Weierstrass theorem, we get
Corollary 4.2.
The $\mathbb{Q}$-vactor space generated by model functions associated to boundary ideal sheaves is dense in $C(\widetilde{U}^{b})$.
Thus it’s ready to prove our main theorem for the non-archimedean case.
Proof of Theorem 3.7.
Let $\bar{D}=(D,g_{D})\in\widehat{\mathrm{Div}}(U^{\mathrm{an}})_{\mathrm{eqv}}$. To show that it comes from $\mathrm{Div}(U/\mathscr{O}_{k})$, it suffices to assume that $D=0$ and show that any continuous norm-equivalent function on $X^{\mathrm{an}}$ is a Green function of some boundary adelic divisors in $\mathrm{Div}(X/\mathscr{O}_{k})$.
Fix a projective model $X$ such that $D_{0}=X\backslash U$ be a divisor. Then $h_{D}=g_{D}/g_{D_{0}}$ is a continuous norm-invariant function on $U^{b}$. By the above proposition, $h_{D}$ can be approximated by a series of model functions $h_{E_{i}}$, where $E_{i}\in\mathrm{Div}(U)_{\mathrm{mod}}$ is a model boundary adelic divisor.
We claim that the sequence $E_{i}$ is a Cauchy sequence. Since $h_{E_{i}}$ is a Cauchy sequence in $C(\widetilde{U}^{b})$, we see that there exists $\varepsilon_{i}\in\mathbb{Q}_{>0}$, such that for all $j>i$
$|h_{E_{i}}(x)-h_{E_{j}}(x)|\leqslant\varepsilon_{i}$
Thus we have
$-\varepsilon_{i}g_{D_{0}}(x)\leqslant g_{E_{i}}(x)-g_{E_{j}}(x)\leqslant\varepsilon_{i}g_{D_{0}}(x)$
for all $j>i$ and all $x\in U^{b}$. Note that all these functions are zero in $\widehat{U}^{\beth}$, hence it holds on $U^{\mathrm{an}}$.
Use [YZ, Lemma 3.3.3], possibly after a normalization, we get inequalities
$-\varepsilon_{i}D_{0}\leqslant E_{i}-E_{j}\leqslant\varepsilon_{i}D_{0}$
in terms of effectivities. Thus $E_{i}$ converges to an adelic divisor $E$.
Finally, we show that the Green function of $E$ is just $g_{D}$. On $U^{b}$, this is done by our choice of $\{E_{i}\}$; on $\widehat{U}^{\beth}$, any norm-equivalent orbit either consists only one point or has closure $[0,\infty]$ in $\widehat{U}^{\beth}$, and in both cases the norm-equivariant condition implies that $g_{D}$ is zero. Since $g_{E_{i}}$ are also zero on $\widehat{U}^{\beth}$, we get that $g_{E_{i}}$ converges to $g_{D}$, thus $g_{E}=g_{D}$, and $(D,g_{D})$ is the analytification of the adelic divisor $D$. ∎
Archimedean case and the global case
Next we prove the archimedean case. Let $U$ be a quasi-projective variety over $\mathbb{C}$, and $X$ is a fixed projective model of $U$ with boundary divisor $D_{0}=X\backslash U$, and let $\|\cdot\|_{\mathrm{Ban}}$ be the Banach norm in §2.2.
Again, we may define the space of model functions on $U^{b}$. The problem is that model functions themselves are not closed under taking minimum. We use Yuan–Zhang’s results instead.
First, we give a lemma, which works for both archimedean and non-archimedean local fields. Let $K$ be a local field with valuation ||$\cdot$|| ⋅ |, let $|\cdot|_{0}$ be the trivial norm on $K$, and let $|\cdot|_{\mathrm{Ban}}=\max\{|\cdot|,|\cdot|_{0}\}$. Then $\mathcal{M}(K,|\cdot|_{\mathrm{Ban}})$ is homeomorphic to $[0,1]$.
Let $U$ be a quasi-projective variety over $K$, and consider the structure morphism $\pi:U^{\mathrm{an}}\to\mathcal{M}(K,|\cdot|_{\mathrm{Ban}})\cong[0,1]$. Let $U_{t}^{\mathrm{an}}$ be the fiber over $t$.
Lemma 4.3.
Let $g\in C(U^{\mathrm{an}}_{1})$ be any continuous function. Extend $g$ to $U^{\mathrm{an}}_{(0,1]}$ norm-equivariantly. Then $g$ extends continuously to a continuous function $\tilde{g}\in C(U^{\mathrm{an}})$ with $\tilde{g}|_{U_{0}^{\mathrm{an}}}=0$ if and only if $g$ grows as $o(g_{D_{0}})$ on $U^{\mathrm{an}}_{1}$ along the boundary.
Proof.
Consider the norm-invariant function $g/g_{D_{0}}$, viewed as a continuous function on $U_{1}^{\mathrm{an}}$. Note that $U^{\mathrm{an}}_{1}$ is also an open subset of $\widetilde{U}^{b}$.
Now $X^{\mathrm{an}}_{1}$ and $\widetilde{U}^{b}$ are two compactification of $U_{1}^{\mathrm{an}}$, and the two conditions are both equivalent to $g/g_{D_{0}}$ converges to zero when $x$ approaches the boundary. ∎
Fix a projective model $X_{0}$ of $U$ with boundary divisor $D_{0}=X_{0}\backslash U$. We also fixed a positive Green function $g_{\bar{D}_{0}}$. Let $X$ be any projective model of $U$, and let $\bar{D}=(D,g_{D})\in\widehat{\mathrm{Div}}(X)$ with $D|_{U}=0$. Then $\bar{D}$ induces a norm-equivariant continuous function $\tilde{g}_{\bar{D}}$ on $U^{\mathrm{an}}$. Let $h_{\bar{D}}=g_{\bar{D}}/g_{\bar{D}_{0}}$ be the norm-invariant function on $U^{b}$, which we viewed as a function on $\widetilde{U}^{b}$. Such functions are called model functions.
Proposition 4.4.
The $\mathbb{Q}$-linear combination of model functions are dense in $C(\widetilde{U}^{b})$.
Proof.
First, we show that model functions separate points. Let $x,y\in\widetilde{U}^{b}$ be two different points. If one of them is archimedean, we can only change the Green function on a small neighborhood in the archimedean part leaving all other place fixed. Otherwise, they are both non-archimedean, and we can use the results for non-archimedean case.
However, the model functions themselves are not stable under taking minimum. We need to check that if $h_{1},h_{2}$ are two model functions, then $\min\{h_{1},h_{2}\}$ is contained in the closure of the space of model functions.
Suppose that $h_{i}$ is the model function of $(D_{i},g_{i})$. Let $\widetilde{X}_{\eta,0}$ be the non-archimedean part of $\widetilde{U}^{b}$. Then $\min\{h_{1},h_{2}\}|_{\widetilde{X}_{\eta,0}}$ is a model function of some boundary divisor $E$. Take any Green function $g_{E}$ of $E$. Now we consider the difference
$\min\{h_{1},h_{2}\}-h_{E}$
It is identically zero on the non-archimedean part. By [YZ, Thm. 3.6.4], it is the Green function of some adelic divisor, which is the limit of a sequence of model functions.
Now by Stone-Weierstrass theorem, we get the density result. ∎
The rest part of the theorem is exactly the same as the non-archimedean case.
The case of $\mathbb{Z}$ is proved the same as the archimedean case.
5 Detailed Structure of $\widetilde{U}^{b}$
We give a more explicit description of $\widetilde{U}^{b}$ when $k=\mathbb{C}$ with either the trivial norm $|\cdot|_{0}$ or the hybrid semi-norm $|\cdot|_{\mathrm{Ban}}=\max\{|\cdot|,|\cdot|_{0}\}$.
If $U$ is quasi-projective variety over $\mathbb{C}$, let $U^{\mathrm{an}}=(U/(\mathbb{C},|\cdot|))^{\mathrm{an}}$ be the usual analytification of $U$, which is a complex analytic space. Let $U_{0}^{\mathrm{an}}=(U/(\mathbb{C},|\cdot|_{0}))^{\mathrm{an}}$ be the non-archimedean analytification of $U$ and $U_{h}^{\mathrm{an}}=(U/(\mathbb{C},|\cdot|_{\mathrm{Ban}}))^{\mathrm{an}}$ be the hybrid analytification of $U$. Then $U^{\mathrm{an}}$ and $U^{\mathrm{an}}_{0}$ are both naturally embedded in $U_{h}^{\mathrm{an}}$.
First assume the $k=\mathbb{C}$ with the trivial norm $|\cdot|_{0}$.
Let $U$ be a smooth quasi-projective variety over $\mathbb{C}$. Let $X$ be a smooth projective model of $U$ such that $D=X\backslash U$ is a divisor. Let $\{D_{i}\}_{i\in I}$ be the set of irreducible components in $D$, and let $\mathrm{Div}_{b}(X,U)_{\mathbb{R}}$ be the $\mathbb{R}$-linear space generated by $\{D_{i}\}$. Then we have a evaluation morphism
$\mathrm{ev}_{X}:U^{\mathrm{an}}\to\mathrm{Div}_{b}(X,U)_{\mathbb{R}}^{*}$
mapping a point $x\in U^{\mathrm{an}}$ to the linear functional $E\mapsto g_{E}(x)$. Such a functional is necessarily semi-positive, i.e. if $E$ is an effective divisor, then $g_{E}(x)\geqslant 0$. The evaluation morphism is $\mathbb{R}_{+}$-equivariant, and maps $\mathrm{ev}_{X}(x)=0$ if and only if $x\in U^{\beth}$.
Now assume further that $D$ is a normal crossing divisor and for each $J\subset I$, $\cap_{j\in J}D_{j}$ is either irreducible or empty. Such a model is called an SNC model. Let $\{e_{i}\}_{i\in I}$ be the dual basis of $\{D_{i}\}_{i\in I}$. For each $J\subset I$, if $\cap_{i\in J}D_{i}$ is non-empty, let $\hat{\sigma}_{J}\subset\mathrm{Div}_{b}(X,U)_{\mathbb{R}}^{*}$ be the simplicial cone $\hat{\sigma_{J}}:=\sum\mathbb{R}_{+}e_{j}$. Let $\hat{\Delta}_{X}$ be the union of all cones $\hat{\sigma}_{J}$, see [BFJ].
Our dual complex $\Delta_{X}$ is defined by the quotient
$\Delta_{X}=(\hat{\Delta}_{X}\backslash\{0\})/\mathbb{R}_{+}.$
We may also fix the boundary divisor $D$, and consider the subspace
$\Delta_{X,D}=\hat{\Delta}_{X}\cap\{e\in\mathrm{Div}_{b}(X,U)_{\mathbb{R}}^{*}:\langle D,e\rangle=1\},$
which is canonically homeomorphic to $\Delta_{X}$.
By exactly the same argument as in [BFJ, §3], we have
Theorem 5.1.
Let $X$ be any SNC model of $U$.
(1) Then image of the evaluation map $\mathrm{ev}_{X}:U^{\mathrm{an}}\to\mathrm{Div}_{b}(X,U)_{\mathbb{R}}^{*}$ coincides with $\Delta_{X}$, and it induces a surjective map $\widetilde{U}^{b}\to\Delta_{X}$.
(2) There is a unique continuous map $\mathrm{emb}_{X}:\Delta_{X}\to\widetilde{U}^{b}$, such that
(i) $\mathrm{ev}_{X}\circ\mathrm{emb}_{X}$ is the identity on $\Delta_{X}$.
(ii) for any $s\in\Delta_{X}$, the reduction of $\mathrm{emb}_{X}(s)$ is the generic point $\xi_{J}$ of $\cap_{j\in J}E_{j}$ for the unique subset $J\subset I$ such that $s$ is contained in the relative interior of $\sigma_{J}$.
See also [Thu], [Nic] for a brief introduction.
Similar as many known results in the non-archimedean analytic geometry, we have
Proposition 5.2.
The evaluation maps $\mathrm{ev}_{X}:\widetilde{U}^{b}\to\Delta_{X}$ induce a homeomorphism
$\mathrm{ev}:\widetilde{U}^{b}\cong\lim_{\begin{subarray}{c}\longleftarrow\\ X\ SNC\end{subarray}}\Delta_{X},$
where $X$ ranges over all SNC model of $U$, and $\Delta_{X}$ is the dual complex of the boundary divisor in $X$.
The proof is almost the same as proofs in [BFJ].
By the last section, any continuous equivariant function on $\widetilde{U}^{b}$ will induce an adelic divisor on $U$. We are especially interested in the hose functions which are the pull back of continuous equivariant functions on $\hat{\Delta}_{X}$ for a fixed SNC model $X$. Such adelic divisors and Green functions are called special adelic divisors with respect to $X$.
Now let $X$ be a fixed SNC model of $U$. Let $D_{i}$ be an irreducible components of $D$. Let $g_{D_{i}}(x)$, $g_{D}(x)$ be the Green functions of $D_{i}$.
Proposition 5.3.
The function $g_{D_{i}}(x)$ is the pull-back of the piecewise-linear function on $\hat{\Delta}_{X}$ with value $1$ on the vertex $e_{i}$ and value 00 on all other vertices.
Proof.
By the definition of the evaluation map, we have
$\langle\mathrm{ev}_{X}(x),D_{i}\rangle=g_{D_{i}}(x),$
thus the Green function $g_{D_{i}}(x)$ is just the coordinate function under the dual basis $\{e_{i}\}_{i\in I}$ of $\{D_{i}\}_{i\in I}$, which is piecewise linear. ∎
In particular, the divisor $D_{i}$ is special with respect to $X$.
If $E$ is special, then for $x\in U^{b}$, $g_{E}(x)/g_{D}(x)$ is $\mathbb{R}_{+}$-invariant, and thus there exists a continuous function $h_{E}(s)$ on $\Delta_{X}$, such that $g_{E}(x)/g_{D}(x)=h_{E}(\mathrm{ev}(x))$. Then $h_{D}(s)=1$ for all $s\in\Delta_{X}$, and $h_{D_{i}}$ are all piecewise linear on $\Delta_{X}$. Since $\Delta_{X}$ is a simplicial complex, the functions $h_{D_{i}}(s)$ separate points on $\Delta_{X}$.
If $\mathcal{I}$ is a boundary ideal sheaf, let $E_{\mathcal{I}}$ be the exceptional divisor of the blowing up of $X$ along $\mathcal{I}$. Then $E_{\mathcal{I}}$ is an adelic divisor, and the corresponding Green function is denoted by $g_{\mathcal{I}}$. In this way, we can construct many adelic divisors.
If $\mathcal{I}_{1}$, $\mathcal{I}_{2}$ are two boundary ideal sheaves, then so is $\mathcal{I}_{1}\cdot\mathcal{I}_{2}$ and $\mathcal{I}_{1}+\mathcal{I}_{2}$, and the corresponding Green function satisfies
$g_{\mathcal{I}_{1}\cdot\mathcal{I}_{2}}(x)=g_{\mathcal{I}_{1}}(x)+g_{\mathcal{I}_{2}}(x)$
$g_{\mathcal{I}+\mathcal{I}_{2}}(x)=\min\{g_{\mathcal{I}_{1}}(x),g_{\mathcal{I}_{2}}(x)\}$
Fix an SNC model $X$ of $U$ with boundary divisor $D$. Let $\{D_{i}\}$ be the irreducible components of $D$, and let $\mathcal{I}_{i}$ be the ideal sheaf of $D_{i}$. If $\vec{t}=(t_{1},\ldots,t_{|I|})\in\mathbb{Z}^{|I|}$, define an ideal sheaf $\mathcal{I}^{\vec{t}}:=\mathcal{I}_{1}^{t_{1}}\cdots\mathcal{I}_{|I|}^{t_{|I|}}$.
For each finite set $T\subset\mathbb{Z}^{|I|}_{\geqslant 0}\backslash\{\vec{0}\}$, define $\mathcal{I}_{T}$ to be the sum of all ideal sheaves $\{\mathcal{I}^{t}\}_{t\in T}$, and $E_{T}$ be the corresponding adelic divisor, and $g_{\mathcal{I}_{T}}$ be its Green function. Then $g_{\mathcal{I}_{T}}$ is also pull-back of a piecewise-linear function on $\Delta_{X}$.
Let $\mathrm{Div}(X,U)_{s}$ be the $\mathbb{Q}$-linear combination of such adelic divisors. Similarly as in [Son, §5.1], the Green functions of these adelic divisors are dense in $C(\Delta_{X})$. By taking limit, any special adelic divisors can be obtained by these model special adelic divisor.
Compare these two argument, we get In the author’s earlier work [Son, §5], we defined a particular adelic divisors called special adelic divisors. These divisors are determined by restrictions to all curves passing to the boundary. We Compare these two constructions, we get:
Proposition 5.4.
Let $U$ is a smooth quasi-projective variety with a fixed projective SNC model $X$ with boundary divisor $D$. Let $E\in\widehat{\mathrm{Div}}(U/\mathbb{C})_{b}$ be an adelic divisor. Then $E$ is a special adelic divisor at every point on $X$ if and only if its normalized Green function $g_{E}/g_{D}$ is the pull-back of some continuous function $h_{E}$ on $\Delta(D)$ via $\rho_{X}$. Further, $h_{E}$ is just the skeletal function on each simplex in the sense of [Son, Def. 5.1] .
Now consider the hybrid case $|\cdot|_{\mathrm{Ban}}=\max\{|\cdot|_{0},|\cdot|\}$. Let $U$ be a quasi-projective smooth variety, and $U(\mathbb{C})$ be the usual complex manifold associated to $U$. Then $\widetilde{U}^{\mathrm{an}}$ is a compactification of $U(\mathbb{C})$, with boundary part isomorphic to $\widetilde{U}_{0}^{b}\cong\displaystyle\lim_{\begin{subarray}{c}\longleftarrow\\ X\ SNC\end{subarray}}\Delta_{X}$. In particular, we get a compactification of $U(\mathbb{C})$ with a non-archimedean boundary.
We can compare our compactification with the one in [AN, §3]. Our boundary is purely non-archimedean, and needs all possible models. Their boundary need only one model each time, and it contain both archimedean and non-archimedean information.
6 Global Monge-Ampère Measure
Let $k=\mathbb{Z}$ or a field, and let $U$ be a quasi-projective flat variety over $k$ of dimension $n+1$. Let $\widehat{\mathrm{Pic}}(U)_{\mathrm{int}}$ be the group of integrable adelic line bundles defined by Yuan–Zhang. They also defined a intersection theory of adelic line bundles in [YZ, §4], which is a symmetric multi-linear map on $\widehat{\mathrm{Pic}}(U)_{\mathrm{int}}^{n+1}$. It can be extend as follows.
Proposition 6.1.
The intersection pairing extends uniquely to a multi-linear map
$\widehat{\mathrm{Pic}}(U)\times\widehat{\mathrm{Pic}}(U)_{\mathrm{int}}^{n}\to\mathbb{R},$
such that if $\bar{L}_{1},\ldots,\bar{L}_{n}\in\widehat{\mathrm{Pic}}(U)_{\mathrm{int}}$ are fixed, then the map $\widehat{\mathrm{Div}}(U)\to\mathbb{R}$ given by $\bar{E}\mapsto(\bar{E}\cdot\bar{L}_{1}\cdots\bar{L}_{n})$ is continuous with respect to the boundary norm on $\widehat{\mathrm{Div}}(U)$. In other word, it is allowed to have at most one non-integrable adelic line bundle.
Proof.
Fix $\bar{L}_{1},\ldots,\bar{L}_{n}\in\widehat{\mathrm{Pic}}(U)_{\mathrm{snef}}$, and let $E\in\widehat{\mathrm{Div}}(U)_{\mathrm{mod}}$ be an effective model divisor. Then $\bar{E}$ is also integrable, and we have that
$(\bar{E}\cdot\bar{L}_{1}\cdots\bar{L}_{n})\geqslant 0.$ (1)
Now let $\bar{E}\in\widehat{\mathrm{Div}}(U)$ be an effective adelic divisor. Then $\bar{E}=\lim_{i\to\infty}\bar{E}_{i}$ for effective model adelic divisors $\bar{E}_{i}$. Fix a boundary divisor $\bar{D}_{0}$. Then there exist $\varepsilon_{i}\in\mathbb{Q}_{+}$ converging to 00, such that
$-\varepsilon_{i}\bar{D}_{0}\leqslant\bar{E}_{i}-\bar{E}_{j}\leqslant\varepsilon_{i}\bar{D}_{0}$
in terms of effectivities for all $j>i$.
Apply $(-\cdot\bar{L}_{1}\cdots\bar{L}_{n})$, by inequality (1), we have
$-\varepsilon_{i}(\bar{D}_{0}\cdot\bar{L}_{1}\cdots\bar{L}_{n})\leqslant(\bar{E}_{i}-\bar{E}_{j}\cdot\bar{L}_{1}\cdots\bar{L}_{n})\leqslant\varepsilon_{i}(\bar{D}_{0}\cdot\bar{L}_{1}\cdots\bar{L}_{n})$
Thus the sequence $(\bar{E}_{i}\cdot\bar{L}_{1}\cdots\bar{L}_{n})$ is a Cauchy sequence, and we define
$(\bar{E}\cdot\bar{L}_{1}\cdots\bar{L}_{n})=\lim_{i\to\infty}(\bar{E}_{i}\cdot\bar{L}_{1}\cdots\bar{L}_{n}).$
For general case, note that any adelic line bundles is the difference of two effective line bundles, we get our results. ∎
In particular, let $\widehat{\mathrm{Div}}(U)_{b}$ be the group of adelic divisors with zero integral part, then by Theorem 3.7, we have that $\widehat{\mathrm{Div}}(U)_{b}\cong C(U^{\mathrm{an}})_{\mathrm{eqv}}$, which is further non-canonically isomorphic to $C(\widetilde{U}^{b})$ for the compact Hausdorff space $\widetilde{U}^{b}$ defined in §2.2. We thus get a morphism
$\widehat{\mathrm{Pic}}(U)_{\mathrm{int}}^{n}\to(C(\widetilde{U}^{b}))^{\vee}\cong\mathcal{M}(\widetilde{U}^{b})$
where $\mathcal{M}(U^{b})$ is the Banach space of signed Radon measures on $U^{b}$, and the latter isomorphism is given by the Riesz representation theorem for measures. We call it the global Monge-Ampere measure associated to the integrable adelic line bundles $\bar{L}_{1},\ldots,\bar{L}_{n}$, denoted by $\mathrm{MA}(\bar{L}_{1},\ldots,\bar{L}_{n})$, and the intersection pairing is also a integration
$(\bar{E}\cdot\bar{L}_{1}\cdots\bar{L}_{n})=\int_{\widetilde{X}}h_{\bar{E}}(x)\ \mathrm{MA}(\bar{L}_{1},\ldots,\bar{L}_{n}),$
where $h_{\bar{E}}=g_{\bar{E}}/g_{\bar{D}_{0}}$ for the fixed boundary divisor $D_{0}$, and $\mathrm{MA}(\bar{L}_{1},\ldots,\bar{L}_{n})$ is also computed via $D_{0}$.
If all $\bar{L}_{i}$ are equal, the measure is denoted by $\mathrm{MA}(\bar{L})$.
To emphasize the independence of $\bar{D}_{0}$, we introduce the following notation.
$\int_{U^{b}}g(x)\mathrm{MA}(\bar{L}_{1},\ldots,\bar{L}_{n}):=\int_{\widetilde{U}^{b}}\frac{g(x)}{g_{\bar{D}_{0}}(x)}\mathrm{MA}(\bar{L}_{1},\ldots,\bar{L}_{n}),$
for any norm-equivariant measurable function on $U^{b}$, where the measure $\mathrm{MA}(\bar{L}_{1},\ldots,\bar{L}_{n})$ on $\widetilde{U}^{b}$ is computed via the same boundary divisor $\bar{D}_{0}$.
Question
Solve the equation $\mathrm{MA}(\bar{L})=\mu$ for a given semi-positive measure $\mu$ and a given line bundle $L\in\mathrm{Pic}(U)_{\mathrm{snef}}$.
We have the following naive conditions for existence and uniqueness of the Monge-Ampère equation.
Definition 6.2.
Let $\bar{E}\in\widehat{\mathrm{Div}}(U)_{b}$ be an integrable adelic divisor. Then $E$ is called vertically-numerically-trivial if
$(\mathscr{O}(\bar{D})\cdot\mathscr{O}(\bar{E})\cdot\bar{L}_{1}\cdots\bar{L}_{n-1})=0$
for all $\bar{D}\in\widehat{\mathrm{Div}}(U)_{b}$ and all integrable adelic line bundles $\bar{L}_{i}$.
We fix a model line bundle $(X,\bar{L}_{0})$ for $(U,L)$. Then any other line bundle $\bar{L}$ with $\bar{L}|_{U}\cong L$ can be written as a sum of $L_{0}$ with a boundary divisor.
Now consider the equation
$\mathrm{MA}(\bar{L})=\mu,$ (2)
which is equivalent to solve the equation $\mathrm{MA}(\bar{L}_{0}+f(x))=\mu$ for integrable norm-equivariant function $f(x)\in C(U^{b})_{\mathrm{eqv}}$.
Proposition 6.3.
(1) If the equation (2) has a solution, then $\bar{L}_{0}$ and $\mu$ satisfy the conditions
[…]
for all vertically-numerically-trivial $\bar{D}\in\widehat{\mathrm{Div}}(U)_{b}$.
(2) If $\bar{L}_{1}$ is a solution, then so is $\bar{L}_{1}+\mathscr{O}(\bar{E})$ for all vertically-numerically-trivial adelic divisor $\bar{E}\in\widehat{\mathrm{Div}}(U)_{b}$.
The proofs are easy. We call them the naive conditions for the Monge-Ampère equation. It would be natural to ask whether these conditions are sufficient under some positivity conditions of $L$ and $\mu$.