### v2

# Norm-equivariant metrized divisors on Berkovich spaces

Abstract
Yuan and Zhang introduced adelic divisors over a quasi-projective varieties in [YZ]. They showed that such divisors will induce norm-equivariant metrized divisors on corresponding Berkovich space. We will show that all norm-equivariant metrized divisors can be obtained this way, and then give some application on the analytic aspect of Yuan–Zhang’s adelic divisors.
1 Introduction
Let $k=\mathbb{Z}$ or a field equipped with a Banach norm $|\cdot|_{\mathrm{Ban}}$. Let $U$ be a quasi-projective integral flat scheme over $k$. Then we obtain a Berkovich analytic space $U^{\mathrm{an}}$ as in [Ber09] containing all possible valuations of residue fields of points in $U$ whose restriction to $k$ is bounded by $|\cdot|_{\mathrm{Ban}}$. It is a Hausdorff, locally compact topological space.
A metrized divisor $\bar{D}=(D,g_{D})$ on $U^{\mathrm{an}}$ consists of a Cartier divisor $D\in\mathrm{Div}(U)$ and a continuous Green function
$g_{D}:U^{\mathrm{an}}\backslash|D|^{\mathrm{an}}\to\mathbb{R},$
such that if $D$ is locally defined by a function $f\in K(U)^{*}$, then $g_{D}(x)+\log|f(x)|$ locally extends continuously to $|D|^{\mathrm{an}}$. The metric is called norm-equivariant if for any $x,y\in U^{\mathrm{an}}\backslash|D|^{\mathrm{an}}$ satisfying $|\cdot|_{x}=|\cdot|_{y}^{t}$ for some $t\in(0,\infty)$, then $g_{D}(x)=tg_{D}(y)$.
Let $\widehat{\mathrm{Div}}(U^{\mathrm{an}})_{\mathrm{eqv}}$ be the group of norm-equivariant metrized divisors on $U^{\mathrm{an}}$.
Recently, Yuan and Zhang introduced adelic divisors on quasi-projective varieties in [YZ]. Let $k=\mathbb{Z}$ or a field, and let $U$ be a quasi-projective flat integral scheme over $k$. Let $\widehat{\mathrm{Div}}(U/k)$ be the group of adelic divisors on $U$ defined in [YZ, §2]. It is a reasonable limit of a sequence of divisors on projective models. Roughly speaking, an adelic divisor $\bar{D}\in\widehat{\mathrm{Div}}(U/k)$ is represented by a Cauchy sequence $\bar{D}_{i}\in\widehat{\mathrm{Div}}(X_{i})_{\mathbb{Q}}$ on different projective models $X_{i}$ of $U$, such that all restrictions $\bar{D}_{i}|_{U}$ are equal and it satisfies a Cauchy condition on the boundary. See §3 for more precise definitions.
Equipped $k$ with either absolute norm(when $k=\mathbb{Z}$) or the trivial norm, they defined a Berkovich analytification map of adelic line bundles in [YZ, §3], and obtain a canonical injection $\widehat{\mathrm{Div}}(U/k)\to\widehat{\mathrm{Div}}(U^{\mathrm{an}})_{\mathrm{eqv}}.$
In our paper, we will show that the above analytification maps are in fact bijective.
Theorem 1.1.
(Theorem 3.7) Let $k$ be either $\mathbb{Z}$ or a field. Let $U$ be a quasi-projective flat integral scheme over $k$. Then the above canonical map
$\widehat{\mathrm{Div}}(U/k)\to\widehat{\mathrm{Div}}(U^{\mathrm{an}})_{\mathrm{eqv}},$
is bijective.
We can also get similar results for local fields $K$, but in this case our analytification map is slightly different from Yuan–Zhang’s one.
Let $K$ be either a complete non-archimedean field with a non-trivial valuation, or $K=\mathbb{R}$ or $\mathbb{C}$ with the standard absolute norm. The norm on $K$ is denoted by ||$\cdot$|| ⋅ |.
Let $|\cdot|_{0}$ be the trivial norm on $K$. Define the hybrid norm as
$|\cdot|_{\mathrm{hyb}}=\max\{|\cdot|,|\cdot|_{0}\}.$
Let $U$ be a quasi-projective variety over $K$. If $K$ is non-archimedean, let $\mathscr{O}_{K}$ be the valuation ring of $K$; if $K$ is archimedean, let $\mathscr{O}_{K}=K$. Then Yuan–Zhang defined the group of local adelic divisors $\widehat{\mathrm{Div}}(U/\mathscr{O}_{K})$. Let $U^{\mathrm{an}}=(U/(K,|\cdot|_{\mathrm{hyb}}))^{\mathrm{an}}$ be the Berkovich space of $U$ over $(K,|\cdot|_{\mathrm{hyb}})$. The local version of Theorem 1.1 is the following.
Theorem 1.2.
(Theorem 3.8) The analytification map
$\widehat{\mathrm{Div}}(U/\mathscr{O}_{K})\to\widehat{\mathrm{Div}}(U^{\mathrm{an}})_{\mathrm{eqv}}$
is bijective.
The article is divided into three parts. In §2, we study some properties of Berkovich spaces that will be used later. We first introduce some background of Berkovich spaces and then we discuss their topological structures.
Let $k=\mathbb{Z}$ or a field with either the usual norm on $\mathbb{Z}$ or the trivial norm. For a quasi-projective variety $U$ over $k$, we introduce two topological spaces related to $U^{\mathrm{an}}$ in §2. We define the interior part $U^{\beth}\subset U^{\mathrm{an}}$ to be the set of points where the reduction map
$\mathrm{red}:U^{\mathrm{an}}\dashrightarrow U$
is well-defined, and define the boundary part
$U^{b}:=U^{\mathrm{an}}\backslash U^{\beth}.$
Let $\widetilde{U}^{b}:=U^{b}/\!\sim_{\mathrm{eqv}}$ be the quotient space of $U^{b}$ modulo the norm-equivalence relation. The space $U^{\beth}$ is closely related to the analytification of formal algebraic $k$-schemes in [Thu].
We will show that the topological spaces $U^{\beth}$ and $\widetilde{U}^{b}$ are both compact Hausdorff. As a byproduct, we obtain a hybrid compactification of quasi-projective analytic varieties. We will compare it with some other compactifications like [AN, BJ] in §2.2. In the case of complex manifolds our definition is essentially the same as in [BJ, Definition 4.9] in suitable cases, but our definition still has many advantages.
The second part is the proof of main theorems. We will show that norm-equivariant metrized adelic divisors are closely related to continuous functions on the topological space $\widetilde{U}^{b}$ discussed in the first part. Further, we will show that the functions related to model divisors form a dense subspace in $C(\widetilde{U}^{b})$, and we get the desired result.
Finally, we give some applications of the main theorems. First we discuss the results on the toric varieties, and obtain a comparison theorem of toric adelic divisor in §4, special adelic divisors in [Son, §5], and toroidal $b$-divisors in [BB]. Then in §5, we discuss the intersection theory of adelic divisors and define the global Monge-Ampère measure on $\widetilde{U}^{b}$ for integrable adelic line bundles which generalizes the Monge-Ampère measure in the complex case, and the Chambert-Loir measure in the non-archimedean case.
Acknowledgement
I would like to thank my advisor Xinyi Yuan for his valuable advice and encouragement for this paper, and thank Ruoyi Guo for his helpful feedback on earlier version of this paper. I would also thank Zhelun Chen for our discussion during the intercity seminar on Arakelov geometry in Kyoto university about toroidal embedding and toroidal $b$-divisors.
2 Berkovich spaces
2.1 Preliminaries on Berkovich spaces
Let $k$ be a commutative ring with unity $1$ which is complete with respect to a norm $|\cdot|_{\mathrm{Ban}}$. Such a pair $(k,|\cdot|_{\mathrm{Ban}})$ is called a Banach ring.
Affine case
Let $(k,|\cdot|_{\mathrm{Ban}})$ be a Banach ring, and $A$ be a $k$-algebra. The Berkovich space $\mathcal{M}(A)$ is the topological space of multiplicative semi-norms $|\cdot|:A\to\mathbb{R}$ whose restriction to $k$ is bounded by $|\cdot|_{\mathrm{Ban}}$. For a point $x\in\mathcal{M}(A)$, the semi-norm is denoted by $|\cdot|_{x}$, and for $f\in A$, the semi-norm $|f|_{x}$ is also denoted by $|f(x)|$, and thus gives a real-valued function $|f|:\mathcal{M}(A)\to\mathbb{R}$. The topology on $\mathcal{M}(A)$ is defined as the weakest topology such that for each $f\in A$, the function $|f|:\mathcal{M}(A)\to\mathbb{R}$ is continuous.
Two points $|\cdot|_{x},|\cdot|_{y}\in\ M(A)$ are called norm-equivalent if there is a positive real number $t\in\mathbb{R}_{+}$ such that $|\cdot|_{x}=|\cdot|_{y}^{t}$.
We discuss two maps between $\operatorname{Spec}A$ and $\mathcal{M}(A)$.
For each $x\in\mathcal{M}(A)$, the kernel of the semi-norm $|\cdot|_{x}:A\to\mathbb{R}$ is a prime ideal. This induces a continuous contraction map
$\rho:\mathcal{M}(A)\to\operatorname{Spec}A$
sending a point $x$ to $\rho(x)=\ker|\cdot|_{x}$. Then the norm $|\cdot|_{x}$ can be viewed as a norm on the residue field at $\rho(x)$. Let $H_{x}$ be the completion of the field $(\kappa(\rho(x)),|\cdot|_{x})$, called the residue field at $x\in\mathcal{M}(A)$. If $H_{x}$ is non-archimedean, let $R_{x}$ be the valuation ring of $H_{x}$.
The other one is the inclusion map $i:\operatorname{Spec}A\to\mathcal{M}(A)$ sending a point $x\in\operatorname{Spec}A$ to the point $(\kappa(x),|\cdot|_{0})$, where $|\cdot|_{0}$ is the trivial norm on $\kappa(x)$, i.e., $|a|_{0}=1$ for all non-zero $a\in\kappa(x)$. This map is not continuous in general.
If $f:A\to B$ is a $k$-algebra homomorphism, $x\in\mathcal{M}(B)$ corresponding to the multiplicative semi-norm $|\cdot|_{x}$, then $|\cdot|_{x}\circ f:A\to\mathbb{R}$ is a multiplicative semi-norm on $A$, thus induces a continuous map $f^{*}:\mathcal{M}(B)\to\mathcal{M}(A)$.
In our paper, we only care about the topological structure of Berkovich spaces, so we do not discuss the notion of analytic functions.
Scheme case
In general, if $(k,|\cdot|_{\mathrm{Ban}})$ is a Banach ring, and $X$ is a scheme over $k$, then $X$ can be covered by open affine subschemes $\{\operatorname{Spec}A_{j}\}$, and we may glue $\mathcal{M}(A_{j})$ together to get the analytification $X^{\mathrm{an}}$, which is more rigorously written as $(X/(k,|\cdot|_{\mathrm{Ban}}))^{\mathrm{an}}$ if we need to emphasize the norm $|\cdot|_{\mathrm{Ban}}$. If $f:X\to Y$ is a morphism of schemes, then $f$ induces a continuous morphism $f^{\mathrm{an}}:X^{\mathrm{an}}\to Y^{\mathrm{an}}$, thus the analytification map is a functor from the category of schemes over $k$ to the category of topological spaces.
Let $(k,|\cdot|_{\mathrm{Ban}})$ be a noetherian Banach ring and $X$ be a scheme of finite type. By [Ber09, Lemma 1.1, Lemma 1.2], we have:
Proposition 2.1.
(1) The space $X^{\mathrm{an}}$ is locally compact.
(1) If $X$ is separated over $k$, then $X^{\mathrm{an}}$ is Hausdorff.
(2) If $X$ is proper over $k$, then $X^{\mathrm{an}}$ is compact.
We see that the contraction maps $\rho_{j}:\mathcal{M}(A_{j})\to\operatorname{Spec}A_{j}$ glue to a continuous map $\rho:X^{\mathrm{an}}\to X$, and the inclusion maps $i_{j}:\operatorname{Spec}A_{j}\to\mathcal{M}(A_{j})$ glue to $i:X\to X^{\mathrm{an}}$. The inclusion map $i:X\to X^{\mathrm{an}}$ is not continuous in general.
Reduction map
For a separated scheme $X$ over $k$, we have a reduction map, which is only partially defined on a subset of $X^{\mathrm{an}}$. Assume that $k$ is a Banach integral domain, and $X$ is a noetherian separated scheme over $k$. Let $X^{\mathrm{an}}=(X/k)^{\mathrm{an}}$. There is a reduction map
$r:X^{\mathrm{an}}\dashrightarrow X$
which is only partially defined on $X^{\mathrm{an}}$ as follows.
For each non-archimedean point $x\in X^{\mathrm{an}}$ such that the restriction of $|\cdot|_{x}$ to $k$ is bounded by the trivial semi-norm on $k$, the canonical map $\operatorname{Spec}H_{x}\to X\to\operatorname{Spec}k$ extends to $\operatorname{Spec}R_{x}\to\operatorname{Spec}k$. Since $X$ is separated, by the valuative criterion, there is at most one $\varphi:\operatorname{Spec}R_{x}\to X$ such that the diagram
commutes. If it does exist, then we say that the reduction map is defined at $x$, and $r(x)$ is the image of the closed points of $\operatorname{Spec}R_{x}$ in $X$ under the extension $\varphi$.
Definition 2.2.
The interior part of $X^{\mathrm{an}}$, denoted by $X^{\beth}\subset X^{\mathrm{an}}$, is the set of all points in $X^{\mathrm{an}}$ such that the reduction map $r$ is defined at $x$. The boundary part $X^{b}$ is defined as $X^{b}=X^{\mathrm{an}}\backslash X^{\beth}$, and the normalized boundary part is $\widetilde{X}^{b}:=X^{b}/\!\sim_{\mathrm{eqv}}$.
Remark 2.3.
Here we explain the notations. If $k$ is a field with trivial norm, in [Thu, §1], Thuillier defined an analytic space $\mathfrak{X}^{\beth}$ for any locally algebraic formal $k$-scheme $\mathfrak{X}$. Then ${X}^{\beth}$ is the analytification of the formal scheme ${X}$(which is just the scheme $X$ with the discrete topology on the structure sheaf).
Let $U$ be a quasi-projective integral flat scheme over $k$, $X$ be a projective model of $U$ and let $\mathfrak{X}$ be the formal completion of $X$ along $D=X\backslash U$. Then $\mathfrak{X}^{\beth}=U^{b}\cup D^{\mathrm{an}}$, and $U^{b}=\mathfrak{X}_{\eta}$ is equal to the generic fiber of the formal scheme $\mathfrak{X}$ in [Thu]. We choose this notation just because it only depends on $U$, not on the projective model, and we want to emphasize that $U^{\mathrm{an}}=U^{\beth}\cup U^{b}$.
If $k$ is a noetherian Banach integral domain with the trivial norm, and $X$ is proper, then the reduction map is defined on the whole space $X^{\mathrm{an}}$, and the map $r:X^{\mathrm{an}}\to X$ is anti-continuous [Ber90, Corollary 2.4.2], namely, the inverse image of an open subset in $X$ is closed in $X^{\mathrm{an}}$.
Now we state the main theorem we need about the Berkovich spaces.
Assumption 2.4.
In this article, a Banach ring $(k,|\cdot|_{\mathrm{Ban}})$ will be one of the following.
  1. (1)
    Trivial norm case: $k$ is any noetherian integral domain with the trivial norm $|\cdot|_{\mathrm{Ban}}=|\cdot|_{0}$.
  2. (2)
    Hybrid case: $k$ is a field with a complete non-trivial valuation ||$\cdot$|| ⋅ |, and take the hybrid norm $|\cdot|_{\mathrm{Ban}}=|\cdot|_{\mathrm{hyb}}=\max\{|\cdot|,|\cdot|_{0}\}$.
  3. (3)
    Global case: $k=\mathbb{Z}$, and the norm $|\cdot|_{\mathrm{Ban}}$ is the standard absolute norm on $\mathbb{Z}$.
Note that in all cases, the trivial norm $|\cdot|_{0}$ is bounded by the Banach norm $|\cdot|_{\mathrm{Ban}}$
Theorem 2.5.
Let $k$ be as in Assumption 2.4, and $U$ be a quasi-projective integral flat scheme over $k$. Then the topological spaces $U^{\beth}$, $\widetilde{U}^{b}$ are both compact Hausdorff.
Before we give the proof of Theorem 2.5, we give a short introduction of Berkovich spaces over fields with hybrid norms.
2.2 Hybrid Berkovich spaces
Now let $K$ be a complete field with respect to a non-trivial norm $|\cdot|_{1}$, and let $|\cdot|_{0}$ be the trivial norm on $K$. Define the hybrid norm $|\cdot|_{\mathrm{hyb}}:=\max\{|\cdot|_{1},|\cdot|_{0}\}$. Then we can check that $|\cdot|_{t}:=|\cdot|_{1}^{t}$ is a multiplicative norm for each $t\in(0,1]$, and $|\cdot|_{1}^{0}=|\cdot|_{0}$. These norms are all bounded by $|\cdot|_{\mathrm{hyb}}$, and it induces a continuous injection $j:[0,1]\to\mathcal{M}(K,|\cdot|_{\mathrm{hyb}})$.
Proposition 2.6.
The injection $j:[0,1]\to\mathcal{M}(K,|\cdot|_{\mathrm{hyb}})$ is a homeomorphism.
Proof.
Let $|\cdot|_{\lambda}$ be any multiplicative norm on $K$ which is bounded by $|\cdot|_{\mathrm{hyb}}$. We will show that there is an element $t\in[0,1]$ such that $|\cdot|_{\lambda}=|\cdot|_{1}^{t}$.
If $|\cdot|_{\lambda}$ is trivial, then we take $t=0$. Otherwise, there exist $\pi\in K$ such that $|\pi|_{\lambda}>1$. Then $|\pi|_{\mathrm{hyb}}\geqslant|\pi|_{\lambda}>1$, hence $|\pi|_{1}>1$. There exist a unique number $t\in(0,1]$, such that $|\pi|_{\lambda}=|\pi|_{1}^{t}$.
For any non-zero $x\in K$, if $|x|_{t}\leqslant 1$, then $|x|_{\mathrm{hyb}}=1$ and hence $|x|_{\lambda}\leqslant 1$. Therefore, for any rational number $a,b\in\mathbb{Q}$ with $|\pi|_{t}^{a}\leqslant|x|_{t}\leqslant|\pi|_{t}^{b}$, we get $|\pi|_{\lambda}^{a}\leqslant|x|_{\lambda}\leqslant|\pi|_{\lambda}^{b}$. But $|\pi|_{\lambda}=|\pi|_{t}$, hence $|x|_{t}=|x|_{\lambda}$ for all $x\in K$. ∎
To simplify the notations, we write $K_{t}$ (resp. $K_{\mathrm{hyb}}$) to be the Banach field $K$ with the norm $|\cdot|_{t}$ (resp. $|\cdot|_{\mathrm{hyb}}$). In the followings we will identify $\mathcal{M}(K_{\mathrm{hyb}})$ with the closed interval $[0,1]$ via $j$. If some statement works for all $K_{t}$ and $K_{\mathrm{hyb}}$, we will write $K_{\bullet}$ instead.
Now let $U$ be any quasi-projective variety over $K$. Then we have the Berkovich space $(U/K_{\mathrm{hyb}})^{\mathrm{an}}$ with the structure map $\pi:(U/K_{\mathrm{hyb}})^{\mathrm{an}}\to[0,1]$. The central fiber is $\pi^{-1}(0)\cong(U/K_{0})^{\mathrm{an}}$, and the fiber over $1$ is $\pi^{-1}(1)\cong(U/K_{1})^{\mathrm{an}}$. For other fibers, we have the following results.
Proposition 2.7.
We have an homeomorphism $\pi^{-1}((0,1])\cong(U/K_{1})^{\mathrm{an}}\times(0,1]$
The proof of archimedean case is given in [Ber09]. The author implicitly uses the following lemma. The proof of non-archimedean case is the same and even easier.
Lemma 2.8.
[…]
Proof.
We only need to check the triangle inequality. By direct computation, we get:
$\|x+y\|^{n}$ $\leqslant\sum_{i=0}^{n}\left|{n\choose i}\right|^{t}\|x\|^{i}\|y\|^{n-i}$
$\leqslant(n+1)\left(\sum_{i=0}^{n}\left|{n\choose i}\right|\|x\|^{\frac{i}{t}}\|y\|^{\frac{n-i}{t}}\right)^{t}$
$=(n+1)\left(\|x\|^{\frac{1}{t}}+\|y\|^{\frac{1}{t}}\right)^{nt}$
Take $nt$-th roots, and let $n\to\infty$, we get
$\|x+y\|^{\frac{1}{t}}\leqslant\|x\|^{\frac{1}{t}}+\|y\|^{\frac{1}{t}}$
Compactifications
The topological spaces $(U/K_{\bullet})^{\mathrm{an}}$ with any norm on $K$ as above are all Hausdorff and locally compact, but it is not compact unless $U$ is projective over $K$. For many purpose we need to obtain compact spaces, and hence we define various compactifications of these spaces, especially of $(U/K_{1})^{\mathrm{an}}$.
The most common one is just given by a compactification of varieties. Namely, given an open immersion of varieties $U\subset X$ with $X$ projective, we obtain an compactifications $(U/K_{\bullet})^{\mathrm{an}}\to(X/K_{\bullet})^{\mathrm{an}}$. Of course it depends on $X$, so it is not intrinsic. One solution is that we can take all projective compactifications $X$ and consider the inverse limit $\lim_{X\supset U}(X/K_{\bullet})^{\mathrm{an}}$, which is a kind of Zariski-Riemann spaces.
For the analytic space $(U/K_{1})^{\mathrm{an}}$, there is another compactification called hybrid compactification, as in [BJ, §2]. Although they only consider the case of complex manifolds, the essential idea works also for general valuation fields.
In the case of complex manifold, they add some non-archimedean part as the boundary, usually the dual complex of boundary divisor. In [BJ, Definition 4.4], they also consider the inverse limit of all such compactifications. This definition is essentially the same as ours defined later.
In [AN], they define the hybrid space of higher ranks, namely, the boundary itself contains both archimedean part and the non-archimedean part.
Using hybrid norms, we see that $(U/K_{\mathrm{hyb}})^{\mathrm{an}}$ contains a canonical subset homeomorphic to $(U/K_{1})^{\mathrm{an}}\times(0,1]$, and thus by Theorem 2.5, we have the following result.
Corollary 2.9.
The topological space
$\widetilde{U}^{b}:=\left((U/K_{\mathrm{hyb}})^{\mathrm{an}}\backslash U^{\beth}\right)/\!\sim_{\mathrm{eqv}}$
is a compactification of $(U/K_{1})^{\mathrm{an}}$. Its boundary is the topological space $\widetilde{U}_{0}^{b}$ derived from the Berkovich space $(U/K_{0})^{\mathrm{an}}$ over the trivially-normed field.
This compactification works for all complete field $K$ with a non-trivial norm $|\cdot|_{1}$, and all quasi-projective varieties $U$(not necessarily smooth).
We can also get other non-archimedean compactifications by retracting the boundary. Let $U$ be a smooth quasi-projective variety with a smooth projective model $X$, such that $D=X\backslash U$ is a simple normal crossing divisor. Let $\Delta(D)$ is the dual complex of $D$. Thuillier has showed the following results in [Thu, Proposition 4.7]
Theorem 2.10.
There is a canonical subset $\mathcal{S}(U)^{*}\subset U_{0}^{b}$, homeomorphic to $\Delta(D)\times(0,\infty)$, such that the topological space $U_{0}^{b}$ retracts to $\mathcal{S}(U)^{*}$.
It can be shown in the proof that this retraction map is equivariant. Hence we get a continuous surjective map $f:\widetilde{U}_{0}^{b}\to\Delta(D)$.
If we take the pushout of the map $f$ and the canonical injection $\widetilde{U}_{0}^{b}\to\widetilde{U}^{b}$, then we get a compactification of the complex manifold $(U/K_{1})^{\mathrm{an}}$ with boundary $\Delta(D)$. This reintroduces the work in [BJ, §2].
2.3 Proof of Theorem 2.5
First, we show that $U^{\beth}$ is compact Hausdorff in all cases, and then so is $\widetilde{U}^{b}$ case by case.
Proposition 2.11.
Let $k$ be as in Assumption 2.4. Then the interior part $U^{\beth}$ is compact Hausdorff.
Proof.
Let $X$ be a projective model of $U$, and $X^{\mathrm{an}}=(X/(k,|\cdot|_{\mathrm{Ban}}))^{\mathrm{an}}$. Since $X$ is proper over $k$, $X^{\beth}$ is just the set of points in $X^{\mathrm{an}}$ such that $|a|_{x}\leqslant 1$ for all $a\in k$. Equivalently, we have
$X^{\beth}\cong(X/(k,|\cdot|_{0}))^{\mathrm{an}},$
and hence $X^{\beth}$ is compact Hausdorff.
By [Ber90], the reduction map $r:X^{\beth}\to X$ is anti-continuous, i.e., the inverse image of an open subset is closed. Thus $U^{\beth}=r^{-1}(U)$ is a closed subspace of $X^{\beth}$, hence it is compact Hausdorff. ∎
Now we show that $\widetilde{U}^{b}$ is compact Hausdorff case by case.
Trivial norm case
In this subsection we prove Theorem 2.5 in the case when $k$ is any noetherian integral domain with the trivial norm $|\cdot|_{\mathrm{Ban}}=|\cdot|_{0}$.
Let $U$ be a quasi-projective integral flat scheme over $k$. Fix a projective model $X/k$ of $U$, and let $D=X\backslash U$, with the reduced subscheme structure. Then the reduction map $r:X^{\mathrm{an}}\to X$ is defined on the whole space $X^{\mathrm{an}}$.
Proof.
First we show the Hausdorffness of $\widetilde{U}^{b}=U^{b}/\!\sim_{\mathrm{eqv}}$.
Let $x,y\in U^{b}$ be two points which are not norm-equivalent. Then $r(x),r(y)$ are two points in $D$(possibly equal). Since $X$ is projective, there exists an open affine subscheme $\operatorname{Spec}A\subset X$ containing $r(x)$ and $r(y)$. Then $\operatorname{Spec}A$ also contains $\rho(x)$ and $\rho(y)$, and $x,y$ give two norm $|\cdot|_{x},|\cdot|_{y}$ on $A$.
Since $r(x),r(y)\in\operatorname{Spec}A$, we have $|a|_{x}\leqslant 1$, $|a|_{y}\leqslant 1$ for all $a\in A$.
Note that $\rho(y)\notin\overline{\{r(x)\}}$, as $\rho(y)\notin D$ and $r(x)\in D$. Thus we have
$\{a\in A:|a|_{x}<1\}\not\subset\{a\in A:|a|_{y}=0\}.$
By prime avoidance, two prime ideals $\{a\in A:|a|_{x}=0\}$ and $\{a\in A:|a|_{y}=0\}$ cannot cover the ideal $\{a\in A:|a|_{x}<1\}$. Thus there exists $f_{x}\in A$ such that $10<|f_{x}|_{x}<$ 1 and $|f_{x}|_{y}\neq 0$.
Similarly, take $f_{y}\in A$ such that $10<|f_{y}|_{y}<$ 1 and $|f_{y}|_{x}\neq 0$. Let $f=f_{x}f_{y}$. Then we have $10<|f|_{x}<$ 1 and $10<|f|_{y}<$ 1.
Since $x,y$ are not norm-equivalent, switch $x,y$ if necessary, we may find $g\in A$, such that
$\frac{\log|g(x)|}{\log|f(x)|}<\frac{\log|g(y)|}{\log|f(y)|}.$
Take any $a\in\mathbb{R}$ such that
$\frac{\log|g(x)|}{\log|f(x)|}<a<\frac{\log|g(y)|}{\log|f(y)|},$
Define two open subset of $(\operatorname{Spec}A)^{\mathrm{an}}$ as
$V_{1}=\{v\in(\operatorname{Spec}A)^{\mathrm{an}}:f(v)\neq 0,1;\ \frac{\log|g(v)|}{\log|f(v)|}<a\}$
$V_{2}=\{v\in(\operatorname{Spec}A)^{\mathrm{an}}:f(v)\neq 0,1;\ \frac{\log|g(v)|}{\log|f(v)|}>a\}$
Here we assume that $\log 0=-\infty<a$ for all real number $a$. Then $V_{1}$, $V_{2}$ are two disjoint open subsets of $(\operatorname{Spec}A)^{\mathrm{an}}$ which are stable under norm-equivalence relation and separate $x,y$. Thus $\widetilde{U}^{b}$ is Hausdorff.
For the compactness of $\widetilde{U}^{b}$, since $X^{\mathrm{an}}$ is compact Hausdorff, it is a $T_{4}$-topological space. As $U^{\beth}$, $D^{\mathrm{an}}$ are both closed in $X^{\mathrm{an}}$, there exist disjoint open subsets $W_{1}\supset U^{\beth}$ and $W_{2}\supset D^{\mathrm{an}}$. Let $W_{3}=X^{\mathrm{an}}\backslash(W_{1}\cup W_{2})$, then $W_{3}$ is closed in $X^{\mathrm{an}}$, and hence compact.
We show that $W_{3}\to U^{b}\to\widetilde{U}^{b}$ is surjective. Given any $x\in U^{b}$, choose an affine open subset $\operatorname{Spec}A$ containing $r(x)$, then $x$ corresponding to a norm $|\cdot|_{x}$ on $A$. Define the line $L:[0,\infty]\to X^{\mathrm{an}}$ by
$L(t)=\begin{cases}i(\rho(x)),&\text{if }t=0;\\ |\cdot|_{x}^{t},&\text{if }t\in(0,\infty);\\ i(r(x)),&\text{if }t=\infty,\end{cases}$
where $i$ is the inclusion map $X\to X^{\mathrm{an}}$. Then we can check that $L$ is continuous. Since $L(0)\in W_{1}$ ,and $L(\infty)\in W_{2}$, and $[0,\infty]$ is connected, two disjoint open subsets $L^{-1}(W_{1})$ and $L^{-1}(W_{2})$ cannot cover $[0,\infty]$. Thus there exist $y\in W_{3}$, and $y$ is norm-equivalent to $x$. Then $W_{3}\to\widetilde{U}^{b}$ is surjective.
Now $\widetilde{U}^{b}$ is a continuous image of a compact space, and hence it is compact. ∎
Hybrid norm case
Now let $k$ be the field with a non-trivial norm ||$\cdot$|| ⋅ |, and consider the hybrid norm $|\cdot|_{\mathrm{hyb}}$. Let $U$ be a quasi-projective variety over $k$, and $U^{\mathrm{an}}=(U/K_{\mathrm{hyb}})^{\mathrm{an}}$ be the Berkovich space with structure morphism $\pi:U^{\mathrm{an}}\to\mathcal{M}(k,|\cdot|_{\mathrm{hyb}})\cong[0,1]$, and let $U_{t}^{\mathrm{an}}=\pi^{-1}(t)$ be the fiber over $t$.
Proof.
Again we first show Hausdorffness.
Let $x,y$ be two points in $U^{b}$ which are not norm-equivalent. If $\pi(x),\pi(y)$ are both non-zero, it follows that $U_{1}^{\mathrm{an}}$ is open in $U^{b}/\!\sim_{\mathrm{eqv}}$, and $U_{1}^{\mathrm{an}}$ is Hausdorff.
If $\pi(x)\neq 0$ and $\pi(y)=0$, we fix an projective model $X$, then the reduction map is defined at $y$, and $r(y)\in X\backslash U$. Choose an open affine subset $\operatorname{Spec}A\subset X$ containing $\rho(x)$ and $r(y)$. The two prime ideals $\{a\in A:|a|_{x}=0\}$ and $\{a\in A:|a|_{y}=0\}$ cannot cover the ideal $\{a\in A:|a|_{y}<1\}$. Thus there exists $f\in A$ with $|f|_{x}\neq 0$ and $10<|f|_{y}<$ 1. By multiplying a nonzero $z\in k$, we may assume that $|f|_{x}\neq 0,1$ and $|f|_{y}\neq 0,1$ hold simultaneously.
If $\pi(x)=\pi(y)=0$, similarly as the trivial norm case, we can also find $f$ such that $|f|_{x}\neq 0,1$ and $|f|_{y}\neq 0,1$ hold simultaneously. The rest part of the proof are similar as the non-archimedean case.
To show the compactness, choose a projective model $X$ and let $D=X\backslash U$. Let $Y=D^{\mathrm{an}}\cup X_{1}^{\mathrm{an}}$, which is closed subset in $X^{\mathrm{an}}$. Then $U^{\beth},Y$ are both compact. Take disjoint open neighborhood $W_{1}\supset U^{\beth}$, $W_{2}\supset D^{\mathrm{an}}$, and let $W_{3}=X^{\mathrm{an}}\backslash(W_{1}\cup W_{2})$. Then similarly $W_{2}$ maps surjectively onto $\widetilde{U}^{b}$, thus $\widetilde{U}^{b}$ is compact. ∎
Global case
For the case $k=\mathbb{Z}$ with the standard absolute norm ||$\cdot$|| ⋅ |, note that
$\mathcal{M}(\mathbb{Z},|\cdot|_{0})\cong\mathcal{M}(\mathbb{Z},|\cdot|_{0})\cup\mathcal{M}(\mathbb{Q},|\cdot|_{\mathrm{hyb}}),$
so it can be easily deduced by the two cases above.
3 Adelic divisors on Berkovich spaces
The adelic divisors and line bundles on quasi-projective varieties were introduced by Yuan–Zhang in [YZ, §2]. We just briefly review their definitions.
3.1 Adelic divisors
Let $k=\mathbb{Z}$ or a field, and let $X$ be a projective flat variety over $k$.
If $k=\mathbb{Z}$, an arithmetic divisor on $X$ is a pair $(D,g_{D})$, where $D$ is a Cartier divisor on $X$, and $g_{D}$ is a continuous real-valued conjugation-invariant function on $(X\backslash D)(\mathbb{C})$(with respect to the analytic topology), such that if $D$ is locally defined by $f$, then $g_{D}(x)+\log|f(x)|$ extends continuously to $D(\mathbb{C})$. Equivalently, $g_{D}$ induces a conjugation-invariant metric on the line bundle $\mathscr{O}(D)$ over $X(\mathbb{C})$, such that the canonical section $1\in\mathscr{O}(D)$ has the norm $\|1\|(x)=e^{-g_{D}(x)}$. The arithmetic divisor $(D,g_{D})$ is called effective if $D$ is effective and $g_{D}$ is non-negative.
If $k$ is a field, by abuse of notation, an arithmetic divisor on $X$ is just a usual divisor on $X$. The set of arithmetic divisors on $X$ is denoted by $\widehat{\mathrm{Div}}(X)$.
Now let $U$ be a quasi-projective flat variety over $k$, and let $X$ be a projective model of $U$. Define the group of $(\mathbb{Q},\mathbb{Z})$-divisors as
$\widehat{\mathrm{Div}}(X,U):=\widehat{\mathrm{Div}}(X)_{\mathbb{Q}}\times_{\mathrm{Div}(U)_{\mathbb{Q}}}\mathrm{Div}(U).$
Thus a $(\mathbb{Q},\mathbb{Z})$-divisor $(\bar{D},D^{\prime})$ consists of an arithmetic $\mathbb{Q}$-divisor $\bar{D}\in\widehat{\mathrm{Div}}(X)_{\mathbb{Q}}$ and a usual divisor $D^{\prime}\in\mathrm{Div}(U)$ such that their image in $\widehat{\mathrm{Div}}(U)_{\mathbb{Q}}$ are the same. By abuse of notations, we just write $\bar{D}\in\widehat{\mathrm{Div}}(X,U)$, and call $D^{\prime}$ the integral part of $\bar{D}$.
The group of model divisors on $U$ is defined as
$\widehat{\mathrm{Div}}(U)_{\mathrm{mod}}:=\lim_{\begin{subarray}{c}\longrightarrow\\ X\end{subarray}}\widehat{\mathrm{Div}}(X,U),$
where $X$ ranges over all projective model of $U$, and the morphisms are pull-back of divisors.
There is a boundary norm on $\widehat{\mathrm{Div}}(U)_{\mathrm{mod}}$, which more rigorously is an extended norm, i.e. it allows that $|D|=\infty$ for some $D$. Choose a fixed projective model $X$ such that $Z=X\backslash U$ is a divisor. If $k=\mathbb{Z}$, we further choose a positive Green function $g_{Z}$, and let $\bar{Z}=(Z,g_{Z})$. Let $\bar{D}\in\widehat{\mathrm{Div}}(U)_{\mathrm{mod}}$. The boundary norm $\|\cdot\|_{\bar{Z}}$ is defined as
$\|\bar{D}\|_{\bar{Z}}=\inf\{\varepsilon\in\mathbb{Q}_{>0}:-\varepsilon\bar{Z}\leqslant\bar{D}\leqslant\varepsilon\bar{Z}\},$
where the partial order is given by effectivity of divisors. If the integral part of $\bar{D}$ is non-zero, such an $\varepsilon$ does not exist, and we assume that $\inf\emptyset=\infty$, thus $\|\bar{D}\|_{\bar{\mathbb{Z}}}=\infty$ in this case.
Different boundary divisors give different boundary norms on $\widehat{\mathrm{Div}}(U)_{\mathrm{mod}}$, but they all induce the same boundary topology. The group of adelic divisor $\widehat{\mathrm{Div}}(U/k)$ is defined as the completion of $\widehat{\mathrm{Div}}(U)_{\mathrm{mod}}$ with respect to any boundary norm.
Let $\bar{D}\in\widehat{\mathrm{Div}}(U/k)$. Then by definition, $\bar{D}$ is represented by a sequence $\bar{D}_{i}\in\widehat{\mathrm{Div}}(X_{i})$, where $X_{i}$ are projective models of $U$. We may assume that if $j>i$, then there is a morphism $\pi_{ji}:X_{j}\to X_{i}$ compatible with open immersions $U\subset X_{i}$, and $(X_{0},\bar{Z}=\bar{D}_{0})$ is a boundary divisor defining the boundary norm. The sequence $\bar{D}_{i}$ satisfies the Cauchy condition that there is a sequence $\epsilon_{i}\in\mathbb{Q}_{+}$ convergent to 00 such that
$\epsilon_{i}\pi_{j0}^{*}\bar{Z}\leqslant\pi_{ji}^{*}\bar{D}_{i}-\bar{D}_{j}\leqslant\varepsilon_{i}\pi_{j0}^{*}\bar{Z}$
for all $j>i$. In particular, $D_{i}|_{U}$ are the same for all $i$, and is denoted by $D|_{U}$, called the integral part of $\bar{D}$.
Let $\widehat{\mathrm{Div}}(U/k)_{b}\subset\widehat{\mathrm{Div}}(U/k)$ be the space of adelic divisors whose integral part is zero. These divisors are also called boundary divisors sometimes, but it is different with the one defining the boundary norm. Hopefully it won’t cause confusions.
3.2 Analytification of Adelic Divisors
Let $k$ be a commutative Banach ring which is also an integral domain, and $X$ is an integral scheme over $k$. An metrized divisor (which is called arithmetic divisor in [YZ, §3.3]) $\bar{D}$ on $X^{\mathrm{an}}$ is a pair $(D,g_{D})$, where $D$ is a usual divisor on $X$, and $g_{D}$ is a continuous function on $X^{\mathrm{an}}\backslash D^{\mathrm{an}}$, such that if $D$ is locally defined by a function $f$, then $g_{D}(x)+\log|f(x)|$ extends continuously to $D^{\mathrm{an}}$.
The metrized divisor $\bar{D}$ is called norm-equivariant if for any $x_{1},x_{2}\in X^{\mathrm{an}}\backslash D^{\mathrm{an}}$ satisfying $|\cdot|_{x_{1}}=|\cdot|^{t}_{x_{2}}$ for some $0<t<\infty$, we have that $g(x_{1})=tg(x_{2})$.
Let $\widehat{\mathrm{Div}}(X^{\mathrm{an}})_{\mathrm{eqv}}$ be the group of norm-equivariant arithmetic divisors. Yuan–Zhang defined an analytification map of adelic divisors.
Let $k$ be either $\mathbb{Z}$ or a field. If $k=\mathbb{Z}$, then $|\cdot|_{\mathrm{Ban}}$ is the usual absolute norm on $\mathbb{Z}$. If $k$ is a field, then $|\cdot|_{\mathrm{Ban}}$ is the trivial norm on $k$.
Proposition 3.1.
[YZ, Prop. 3.3.1] Let $X$ be a flat and quasi-projective integral scheme over $k$, and let $X^{\mathrm{an}}=(X/(k,|\cdot|_{\mathrm{Ban}}))^{\mathrm{an}}$. Then there is a canonical injection map
$\widehat{\mathrm{Div}}(X/k)\to\widehat{\mathrm{Div}}(X^{\mathrm{an}})_{\mathrm{eqv}}.$
We briefly discuss their construction. First assume that $X$ is projective. If $k$ is a field, then there is a reduction map $r:X^{\mathrm{an}}\to X$. Let $D$ be a usual divisor on $X$, and let $x\in X^{\mathrm{an}}$. Take any open affine neighborhood $U=\operatorname{Spec}A$ of $r(x)$, such that $D|_{U}$ is locally defined by $f\in k(U)^{*}$. Define $\tilde{g}(x)=-\log|f(x)|$. It is independent of the choice $(U,f)$. Then $(D,\tilde{g})$ is the analytification of $D\in\widehat{\mathrm{Div}}(X/k)$.
If $k=\mathbb{Z}$, let $\bar{D}=(D,g)\in\widehat{\mathrm{Div}}(X)$. The analytification $(D,\tilde{g})$ is defined as follows. The restriction of $\tilde{g}$ to the non-archimedean part of $X^{\mathrm{an}}$ are defined in the same way as above. The archimedean part of $X^{\mathrm{an}}$ is homeomorphic to $X(\mathbb{C})\times(0,1]$ modulo complex conjugaion, thus the restriction of $\tilde{g}$ to the archimedean part is uniquely determined by $g$.
Let $X$ be a fixed projective model of $U$, such that $D=X\backslash U$ is a divisor. Let $\bar{D}$ be a boundary divisor defining the boundary norm, which induces a Green function $g_{\bar{D}}$ on $U^{\mathrm{an}}$. By the above construction, we have the following result.
Proposition 3.2.
For any $x\in U^{\mathrm{an}}$, we have that $g_{\bar{D}}(x)\geqslant 0$, and $g_{\bar{D}}(x)=0$ if and only if $x\in U^{\beth}$.
In general, if $U$ is quasi-projective, one may take limits. See [YZ, §3] for details.
Using Green functions on $U^{\mathrm{an}}$, we can get another topological space which is homeomorphic to $\widetilde{U}^{b}$. Roughly, instead of taking quotient, we choose a representative in each norm-equivalent class according the value of Green functions. Let $U$ be quasi-projective variety over $k$, $X$ be a projective model of $U$, and $\bar{D}$ is a boundary divisor defining the boundary norm.
Definition 3.3.
Let $X,U,\bar{D}$ as above. The subspace $\delta_{\bar{D}}(U)$ be the subspace of $U^{\mathrm{an}}$ defined as follows. If $k$ is a field, then
$\delta_{D}(U)=\{x\in U^{\mathrm{an}}:g_{D}(x)=1\}.$
If $k=\mathbb{Z}$, then it has an extra subset on the archimedean part
$\delta_{\bar{D}}(U)=$$\{x\in U^{b}:g_{\bar{D}}(x)=1\}$
$\cup\{x\in U^{b}:\max_{y\sim_{\mathrm{eqv}}x}{g_{\bar{D}}(y)}\leqslant 1,\text{ and }g_{\bar{D}}(x)=\max_{y\sim_{\mathrm{eqv}}x}{g_{\bar{D}}(y)}\}.$
Proposition 3.4.
The space $\delta_{\bar{D}}(U)$ is compact Hausdorff.
Proof.
If $D$ is locally defined by a function $f$, then $g_{\bar{D}}(x)+\log|f(x)|$ extends continuously to $D^{\mathrm{an}}$. Thus $g_{\bar{D}}:U^{\mathrm{an}}\to\mathbb{R}$ extends continuously to $\bar{g}_{\bar{D}}:X^{\mathrm{an}}\to\mathbb{R}\cup\{\infty\}$, and the space
$\{x\in U^{b}:g_{\bar{D}}(x)=1\}=\bar{g}_{\bar{D}}^{-1}(1)$
is compact Hausdorff. The subspace $\{x\in U^{b}:g_{\bar{D}}(x)=\displaystyle\max_{y\sim_{\mathrm{eqv}}x}{g_{\bar{D}}(y)}\leqslant 1\}$ is homeomorphic to the space
$\{x\in U/(\mathbb{C},|\cdot|)^{\mathrm{an}}:g_{\bar{D}}(x)\leqslant 1\},$
which is also compact Hausdorff. ∎
Corollary 3.5.
The composition
$\delta_{\bar{D}}(U)\to U^{b}\to\widetilde{U}^{b}$
is a homeomorphism.
Proof.
It is a continuous bijection from a compact space to a Hausdorff space, hence it is a homeomorphism. ∎
Corollary 3.6.
If $(k,|\cdot|_{\mathrm{Ban}})$ is non-archimedean, then we have homeomorphism
$U^{b}\cong\widetilde{U}^{b}\times\mathbb{R}$
The space $\delta_{\bar{D}}(U)$ is easier to study, but it depends on the choice of the boundary divisor $\bar{D}$. The space $\widetilde{U}^{b}$ is intrinsic. However, a Green function $g_{\bar{D}}$ does not induce any reasonable function on $\widetilde{U}^{b}$. We can only consider the function $g_{\bar{D}_{1}}/g_{\bar{D}_{2}}$, so we still need to choose a boundary divisor.
Our main theorem is as follows.
Theorem 3.7.
Let $k$ be either $\mathbb{Z}$ or a field. Let $U$ be a flat quasi-projective integral scheme over $k$. Then the above canonical maps
$\widehat{\mathrm{Div}}(U/k)\to\widehat{\mathrm{Div}}(U^{\mathrm{an}})_{\mathrm{eqv}},$
$\widehat{\mathrm{Div}}(U/k)_{b}\to C(U^{\mathrm{an}})_{\mathrm{eqv}}$
are isomorphisms.
Let $K$ be a field complete with respect to a non-trivial discrete valuation ||$\cdot$|| ⋅ |. If $K$ is non-archimedean, let $O_{K}$ be the valuation ring of $K$. If $K$ is archimedean, let $O_{K}=K$. Let $U$ be a quasi-projective variety over $K$.
Yuan and Zhang also define the group of adelic divisors $\widehat{\mathrm{Div}}(U/\mathscr{O}_{k})$ in the this local case, and define injections $\widehat{\mathrm{Div}}(U/\mathscr{O}_{K})\to\widehat{\mathrm{Div}}((U/(K,|\cdot|))^{\mathrm{an}})$. However, to get a similar results as in the global case, we need a different analytification using hybrid norms.
Let $(K,|\cdot|)$ as above, and let $|\cdot|_{0}$ be the trivial norm on $K$. Define the hybrid norm $|\cdot|_{\mathrm{hyb}}=\max\{|\cdot|,|\cdot|_{0}\}$ on $K$, and let $U^{\mathrm{an}}=(U/(K,|\cdot|_{\mathrm{hyb}}))^{\mathrm{an}}$. Similarly as in the global case, there is a natural analytification map $\widehat{\mathrm{Div}}(U/\mathscr{O}_{K})\to\widehat{\mathrm{Div}}(U^{\mathrm{an}})_{\mathrm{eqv}}$.
Theorem 3.8.
Notations are as above. Then the natural map $\widehat{\mathrm{Div}}(U/\mathscr{O}_{K})\to\widehat{\mathrm{Div}}(U^{\mathrm{an}})_{\mathrm{eqv}}$ is an isomorphism.
We give a direct corollary here. Let $U$ be a quasi-projective variety over a complete field $K$ with a non-trivial norm ||$\cdot$|| ⋅ |. Fix a projective model $X$ with $D$ a divisor, and let $g_{D}$ be a positive continuous Green function of $D$ on $U_{1}^{\mathrm{an}}=(U/(K,|\cdot|))^{\mathrm{an}}$. Let $g:U_{1}^{\mathrm{an}}\to\mathbb{R}$ be any continuous function.
Corollary 3.9.
The function $g$ is a Green function of some adelic divisor if and only if the quotient $g/g_{D}:U^{\mathrm{an}}\to\mathbb{R}$ extends continuously to $\widetilde{U}^{b}$, where $\widetilde{U}^{b}$ is viewed as a compactification of $U_{1}^{\mathrm{an}}$.
3.3 Proof of the Main Theorem
We first prove the case that $k=\mathbb{Z}$ or a field with the trivial norm. Then we prove the hybrid case and the global case. It suffices to prove the case when the integral part is zero.
Trivial norm case
First we consider the trivial norm case, where $k=\mathbb{Z}$ or a field with the trivial norm $|\cdot|_{0}$.
Let $X$ be a fixed projective model of $U/k$, such that $D=X\backslash U$ is a divisor. Then each ideal sheaf $\mathcal{I}\subset\mathscr{O}_{X}$ defines a closed subscheme $Y\subset X$.
Now suppose that $|Y|\subset|D|$. Such ideal sheaves are called boundary ideal sheaves. Then $\mathrm{Bl}_{Y}X$ ia also a projective model of $U$, and the exceptional divisor $E_{Y}$ is a model divisor in $\mathrm{Div}(U)_{\mathrm{mod}}$. The analytification map gives an arithmetic divisor $(0,g_{Y})$ on $U^{\mathrm{an}}$ with integral part zero, and a continuous function $g_{Y}(x)$ on $U^{\mathrm{an}}$.
The function $g_{Y}$ has the form
$g_{Y}(x)=\min\{-\log|a(x)|:a\in\mathcal{I}_{r(x)}\}.$
It is identically zero on $U^{\beth}$, and is always norm-equivariant and non-negative. If further that $Y=D$, we see that $g_{D}(x)$ is strictly positive on $U^{b}$.
Since $g_{Y}(x)$ and $g_{D}(x)$ are both norm-equivariant, the quotient $h_{Y}(x)=\frac{g_{Y}(x)}{g_{D}(x)}$ is a norm-invariant continuous function on $U^{b}$, thus it is the pull-back of a continuous function on $\widetilde{U}^{b}$, which is still denoted by $h_{Y}$. Such a function is called a model function on $\widetilde{U}^{b}$ associated to the closed subscheme $Y$. If $Y$ is defined by the boundary ideal sheaf $\mathcal{I}$, we may also write $g_{\mathcal{I}}=g_{Y}$ and $h_{\mathcal{I}}=h_{Y}$.
The constant function $1=h_{D}$ is a model function. Since we have
$h_{\mathcal{I}_{1}\mathcal{I}_{2}}=h_{\mathcal{I}_{1}}+h_{\mathcal{I}_{2}}$
and
$h_{\mathcal{I}_{1}+\mathcal{I}_{2}}=\min\{h_{\mathcal{I}_{1}},h_{\mathcal{I}_{2}}\},$
the space of model functions are closed under addition and taking minimum. To show that the $\mathbb{Q}$-linear combination of model functions associated to the boundary ideal sheaf is dense in the Banach space of continuous functions on $\widetilde{U}^{b}$, we need to show that model functions separate points.
Proposition 3.10.
Model functions on $\widetilde{U}^{b}$ separate points.
Proof.
The proof here is almost the same as in [BFJ, Prop. 2.2]
Let $x,y\in U^{b}$ are two points which are not norm-equivalent. Choose a suitable $x^{\prime}$ which is norm-equivalent to $x$ if necessary, we may assume that $g_{D}(x)=g_{D}(y)$, and still $x\neq y$.
If $r(x)\neq r(y)$, we may assume that $r(y)\notin\overline{r(x)}$. Let $Y=\overline{r(x)}$. Then $h_{Y}(x)>0$ and $h_{Y}(y)=0$, thus $h_{Y}$ separates $x,y$.
Now assume that $r(x)=r(y)$. Take an affine open neighborhood $\operatorname{Spec}A$ of $r(x)$. Since $x\neq y$, there exists $f\in A$ such that $|f(x)|\neq|f(y)|$. Consider the ideal sheaf $\mathscr{O}_{\operatorname{Spec}A}\cdot f$, which can be extended to an ideal sheaf $\mathcal{I}$ on $X$. For each positive integer $m$, the ideal sheaf $\mathcal{I}_{m}=\mathcal{I}+(\mathcal{I}_{D})^{m}$ is a boundary ideal sheaf on $X$, and we have
$h_{\mathcal{I}_{m}}=\min\{\frac{-\log|f|}{g_{D}},m\}$
at $x$ and $y$, hence it separates $x,y$ for $m$ large. ∎
By Boolean ring version of Stone-Weierstrass theorem, we get
Corollary 3.11.
The $\mathbb{Q}$-vactor space generated by model functions associated to boundary ideal sheaves is dense in $C(\widetilde{U}^{b})$.
Thus it’s ready to prove our main theorem for the non-archimedean case.
Proof.
It suffices to show that any continuous norm-equivariant function $g$ on $X^{\mathrm{an}}$ is a Green function of some boundary adelic divisors in $\mathrm{Div}(X/k)$.
Fix a projective model $X$ such that $D_{0}=X\backslash U$ be a divisor. Then $h=g/g_{D_{0}}$ is a continuous norm-invariant function on $U^{b}$. By the above proposition, $h$ can be approximated by a sequence of model functions $h_{E_{i}}$ uniformly on $\widetilde{U}^{b}$, where $E_{i}\in\mathrm{Div}(U)_{\mathrm{mod}}$ is a model boundary adelic divisor.
We claim that the sequence $E_{i}$ is a Cauchy sequence. Since $h_{E_{i}}$ is a Cauchy sequence in $C(\widetilde{U}^{b})$, we see that there exists $\varepsilon_{i}\in\mathbb{Q}_{>0}$, such that for all $j>i$
$|h_{E_{i}}(x)-h_{E_{j}}(x)|\leqslant\varepsilon_{i}$
Thus we have
$-\varepsilon_{i}g_{D_{0}}(x)\leqslant g_{E_{i}}(x)-g_{E_{j}}(x)\leqslant\varepsilon_{i}g_{D_{0}}(x)$
for all $j>i$ and all $x\in U^{b}$. Note that all these functions are zero in $\widehat{U}^{\beth}$, hence it holds on $U^{\mathrm{an}}$.
Use [YZ, Lemma 3.3.3], possibly after a normalization, we get inequalities
$-\varepsilon_{i}D_{0}\leqslant E_{i}-E_{j}\leqslant\varepsilon_{i}D_{0}$
in terms of effectivities. Thus $E_{i}$ converges to an adelic divisor $E$. It is easy to check that the Green function of $E$ is just $g$. ∎
Hybrid norm case and the global case
Next we prove the hybrid norm case. Let $U$ be a quasi-projective variety over $K$, and $X$ is a fixed projective model of $U$ with boundary divisor $D_{0}=X\backslash U$, and let $\|\cdot\|_{\mathrm{Ban}}$ be the hybrid norm.
Again, we may define the space of model functions on $U^{b}$. The problem is that model functions themselves are not closed under taking minimum. We use Yuan–Zhang’s results instead.
First, we give a lemma, which works for both archimedean and non-archimedean local fields. Let $K$ be a local field with valuation ||$\cdot$|| ⋅ |, let $|\cdot|_{0}$ be the trivial norm on $K$, and let $|\cdot|_{\mathrm{hyb}}=\max\{|\cdot|,|\cdot|_{0}\}$. Then $\mathcal{M}(K,|\cdot|_{\mathrm{hyb}})$ is homeomorphic to $[0,1]$.
Let $U$ be a quasi-projective variety over $K$, and consider the structure morphism $\pi:U^{\mathrm{an}}\to\mathcal{M}(K,|\cdot|_{\mathrm{Ban}})\cong[0,1]$. Let $U_{t}^{\mathrm{an}}$ be the fiber over $t$.
Lemma 3.12.
Let $g\in C(U^{\mathrm{an}}_{1})$ be any continuous function. Extend $g$ to $U^{\mathrm{an}}_{(0,1]}$ norm-equivariantly. Then $g$ extends continuously to a continuous function $\tilde{g}\in C(U^{\mathrm{an}})$ with $\tilde{g}|_{U_{0}^{\mathrm{an}}}=0$ if and only if $g$ grows as $o(g_{D_{0}})$ on $U^{\mathrm{an}}_{1}$ along the boundary.
Proof.
Consider the norm-invariant function $g/g_{D_{0}}$, viewed as a continuous function on $U_{1}^{\mathrm{an}}$. Note that $U^{\mathrm{an}}_{1}$ is also an open subset of $\widetilde{U}^{b}$.
Now $X^{\mathrm{an}}_{1}$ and $\widetilde{U}^{b}$ are two compactification of $U_{1}^{\mathrm{an}}$, and the two conditions are both equivalent to $g/g_{D_{0}}$ converges to zero when $x$ approaches the boundary. ∎
Fix a projective model $X_{0}$ of $U$ with boundary divisor $D_{0}=X_{0}\backslash U$. We also fixed a positive Green function $g_{\bar{D}_{0}}$. Let $X$ be any projective model of $U$, and let $\bar{D}=(D,g_{D})\in\widehat{\mathrm{Div}}(X)$ with $D|_{U}=0$. Then $\bar{D}$ induces a norm-equivariant continuous function $\tilde{g}_{\bar{D}}$ on $U^{\mathrm{an}}$. Let $h_{\bar{D}}=g_{\bar{D}}/g_{\bar{D}_{0}}$ be the norm-invariant function on $U^{b}$, which we viewed as a function on $\widetilde{U}^{b}$. Such functions are called model functions.
Proposition 3.13.
The $\mathbb{Q}$-linear combination of model functions are dense in $C(\widetilde{U}^{b})$.
Proof.
First, we show that model functions separate points. Let $x,y\in\widetilde{U}^{b}$ be two different points. If one of them is archimedean, we can only change the Green function on a small neighborhood in the archimedean part leaving all other place fixed. Otherwise, they are both non-archimedean, and we can use the results for non-archimedean case.
However, the model functions themselves are not stable under taking minimum. We need to check that if $h_{1},h_{2}$ are two model functions, then $\min\{h_{1},h_{2}\}$ is contained in the closure of the space of model functions.
Suppose that $h_{i}$ is the model function of $(D_{i},g_{i})$. Let $\widetilde{U}^{b}_{0}$ be the non-archimedean part of $\widetilde{U}^{b}$. Then $\min\{h_{1},h_{2}\}|_{\widetilde{U}^{b}_{0}}$ is a model function of some boundary divisor $E$. Take any Green function $g_{E}$ of $E$. Now we consider the difference
$\min\{h_{1},h_{2}\}-h_{E}$
It is identically zero on the non-archimedean part. By [YZ, Thm. 3.6.4], it is the Green function of some adelic divisor, which is the limit of a sequence of model functions.
Now by Stone-Weierstrass theorem, we get the density result. ∎
The rest part of the theorem is exactly the same as the non-archimedean case, and the case of $\mathbb{Z}$ is deduced by the two cases above.
4 Adelic divisors on toric varieties
In this section we study a class of adelic divisors on toric varieties. Most background we need about toric varieties can be found in the books [CLS, KKMS]. The field $k$ in this section is assumed to be algebraically closed.
4.1 Preliminaries on toric varieties
Let $T\cong(\mathbb{G}m)^{n}$ be a torus over a field $k$. Let $M=\mathrm{Hom}(T,\mathbb{G}m)$ be the character group, and $N=\mathrm{Hom}(\mathbb{G}m,T)$ be the cocharacter group. For each $m\in M$, let $\chi^{m}$ be the corresponding invertible function on $T$. Then $\Gamma(T,\mathscr{O}_{T})$ is generated by $\{\chi^{m}\}_{m\in M}$ as vector space and we have a canonical isomorphism $T\cong\operatorname{Spec}k[M]$.
There is a perfect pairing $M\times N\to\mathbb{Z}$ given by $(a,b)\mapsto\deg a\circ b:\mathbb{G}m\to\mathbb{G}m$. Then it induces the perfect pairing $M_{\mathbb{R}}\times N_{\mathbb{R}}\to\mathbb{R}$. We will always identify $M\cong N^{\vee}$ and $M_{\mathbb{R}}\cong N_{\mathbb{R}}$.
A toric variety $X$ is a variety containing an open subset $X_{0}\cong T$ such that the action $T\times X_{0}\to X_{0}$ extends to an action $T\times X\to X$. There are deep connections between toric varieties and polytopes, polyhedra, and combinatorics. For example, there is a one-to-one correspondence between complete normal toric varieties and a finite rational polyhedral decomposition of $N_{\mathbb{R}}$, see [KKMS, Theorem 6,7,8].
Let $X$ be a complete normal toric variety. Each Cartier divisor supported in the boundary is also a Weil divisor, which is sum of codimension 1 subvarieties contained in $X\backslash X_{0}$. Since $T$ acts on $X$. it must be sum of codimension 1 orbits of $T$. In particular, any boundary Weil divisor and Cartier divisors are $T$-invariant.
Any such a Cartier divisor $D$ will induce a function on $N_{\mathbb{R}}$ called the supporting function in [CLS, §4.] and order function in [KKMS, Theorem 9.]. We will call it supporting function and denote it by $\mathrm{SF}_{D}:N_{\mathbb{R}}\to\mathbb{R}$. Here we do not give the whole definition of the supporting function, but give all properties we need to determine this function.
A point $a\in N$ corresponding to a 1-parameter subgroup $\lambda_{a}:\mathbb{G}m\to X$. Since $X$ is complete, $\lambda_{a}$ extends to a function $\bar{\lambda}_{a}:\mathbb{A}^{1}\to X$. Then the supporting function $\mathrm{SF}_{D}:N_{\mathbb{R}}\to\mathbb{R}$ is the unique continuous conical(meaning that $f(\lambda x)=\lambda f(x)$ for all $\lambda\in\mathbb{R}_{+}$) function such that $\mathrm{SF}_{D}(a)=\mathrm{ord}_{0}\bar{\lambda}_{a}^{*}D$ for all $a\in N$.
4.2 Analytification and tropicalization
In this subsection we study some analytic property of toric varieties. Equipped $k$ with the trivial norm, we get an analytic space $T^{\mathrm{an}}$. Further, there is a tropicalization map
$\mathrm{trop}:T^{\mathrm{an}}\to N_{\mathbb{R}}$
mapping a point $|\cdot|_{x}$ to the linear functional $(\chi^{m}\mapsto-\log|\chi^{m}|_{x})_{m\in M}$. It is $\mathbb{R}_{+}$-equivariant.
If we fix a coordinate $T=\operatorname{Spec}k[T_{1}^{\pm 1},\ldots,T_{n}^{\pm 1}]$, then $M$, $N$ have canonical basis $\{T_{i}\}$ and dual basis $\{T_{i}^{\vee}\}$. Under this basis, the tropicalization map $\mathrm{trop}$ has an explicit form
$\mathrm{trop}:T^{\mathrm{an}}$ $\to\mathbb{R}^{n}$
$|\cdot|_{x}$ $\mapsto(-\log|T_{1}|_{x},\ldots,-\log|T_{n}|_{x}).$
The tropicalization map $\mathrm{trop}:T^{\mathrm{an}}\to N_{\mathbb{R}}$ has a continuous section
$\mathrm{emb}:N_{\mathbb{R}}\to T^{\mathrm{an}}$
sending the linear functional $a:M\to\mathbb{R}$ to the norm
$\left|\sum_{m\in M}a_{m}\chi_{m}\right|=\max_{m\in M}|a_{m}|\mathrm{e}^{-\langle m,a\rangle}.$
Let $q=\mathrm{trop}\circ\mathrm{emb}.$ Then Thuillier showed in [Thu, §2] that $p$ is a continuous retraction of $T^{\mathrm{an}}$ to the its image $q(T^{\mathrm{an}})$. In particular, $q^{2}=q$ and $\mathrm{trop}(x)=\mathrm{trop}(q(x))$.
We are interested in adelic divisors on $T$ that are limit of T-invariant divisors on different toric varieties $X$. First, we give a new description of toric divisors.
Proposition 4.1.
Let $X$ be a complete toric variety, and $D\in\mathrm{Div}(X)$ with $D|_{T}=0$. Let $g_{D}:(X\backslash|D|)^{\mathrm{an}}\to\mathbb{R}$ be Green function of $D$. Then its restriction to $T^{\mathrm{an}}$ is the pull-back of the supporting function $\mathrm{SF}_{D}$ under the tropicalization map $\mathrm{trop}:T^{\mathrm{an}}\to N_{\mathbb{R}}$.
Proof.
We first prove that for any $x,y\in T^{\mathrm{an}}$ with $\mathrm{trop}(x)=\mathrm{trop}(y)$, $g_{D}(x)=g_{D}(y)$. Since $q(x)=q(y)$, we may assume that $y=q(x)$.
Note that $D$ is $T$-invariant, which means that for the two morphism $m,p_{2}:T\times X\to X$, we have $m^{*}D=p_{2}^{*}D$.
Let $g_{D}:(X\backslash|D|)^{\mathrm{an}}\to\mathbb{R}$ be the Green function of $D$. Then the Green function of $m^{*}D$ is just the pull back of $g_{D}$ via the multiplication morphism
$m:(T\times(X\backslash D))^{\mathrm{an}}\to(X\backslash D)^{\mathrm{an}}$
and similarly for $p_{2}^{*}D$. Since $m^{*}D=p_{2}^{*}D$(as usual divisors, not adelic divisors), for each point $z$ in $(T\times X)^{\beth}\backslash(T\times D)^{\beth}$, we must have $g_{D}(m(z))=g_{D}(p_{2}(z))$. Now it follows from Lemma 4.2 below, and $g_{D}$ is the pull-back of some continuous conical function $f$ on $N_{\mathbb{R}}$.
To check that $f=\mathrm{SF}_{D}$, for each $a\in N$, the map $\lambda_{a}:\mathbb{G}m\to T$ induces a map $\lambda_{a}^{\mathrm{trop}}:N_{\mathbb{G}m,\mathbb{R}}\to N_{T,\mathbb{R}}$, which is just the linear map sending the canonical base in $N_{\mathbb{G}m}$ to $a\in N$, and the diagram
commutes. Then it reduces to the case that $T=\mathbb{G}m$, and then by direct computation. ∎
Lemma 4.2.
Take any point $x\in X_{0}^{\mathrm{an}}\cong T^{\mathrm{an}}$, then there is a point $z\in(T\times X_{0})^{\mathrm{an}}$, such that $p_{1}(z)=q(x)\in T^{\beth}$, $p_{2}(z)=x\in X_{0}^{\mathrm{an}}$ and $m(z)=q(x)$.
We will give an rather explicit proof, but the idea is the same as in [Thu].
Proof.
Recall that $q(x)$ is the norm given by the formula
$\left|\sum_{m\in M}c_{m}\chi^{m}\right|_{q(x)}=\max_{m\in M}|c_{m}||\chi^{m}|_{x}$
We check directly that the norm
$\left|\sum_{m,n\in M}c_{mn}\chi^{m}\otimes\chi^{n}\right|_{z}=\max_{m\in M}\left(\left|\sum_{n\in M}c_{mn}\chi^{n}\right|_{x}\right)$
satisfies our conditions. ∎
4.3 Toric adelic divisors
Now we introduce the toric adelic divisors on $T$ and then compare it to toroidal $b$-divisors in [BB] and special adelic divisors in [Son].
Let $f:N_{\mathbb{R}}\to\mathbb{R}$ be any continuous conical function on $N_{\mathbb{R}}$. Since $\mathrm{trop}$ is equivariant, $f\circ\mathrm{trop}$ is an equivariant function on $T^{\mathrm{an}}$, and thus induces an adelic divisor. We will give a detailed description of such adelic divisors.
Definition 4.3.
A toric adelic divisor $D$ on a torus $T$ is an adelic divisor such that $D|_{T}=0$ and its Green function $g_{D}$ is the pull-back of a continuous conic function on $N_{\mathbb{R}}$ under the tropicalization map.
Similarly as the model case, if an adelic divisor $D$ is toric, then the function $f$ has an explicit description.
Proposition 4.4.
If $D$ is a toric adelic divisor, then for each $a\in N$ corresponding to the 1-dimensional parameter subgroup $\lambda_{a}:\mathbb{G}m\to T$, we have
$\lambda_{a}^{*}D=f(a)[0]+f(-a)[\infty],$
where $[0]$ means the adelic divisor $[0]\in\mathrm{Div}(\mathbf{P}^{1})\to\widehat{\mathrm{Div}}(\mathbb{G}_{m})$, and similarly for []delimited-[][$\infty$][ ∞ ].
The proof of it is th same as the model case. Hence the function $f$ is also called the supporting function, and is denoted by $\mathrm{SF}_{D}:N_{\mathbb{R}}\to\mathbb{R}$.
Now we are ready to prove the main theorem.
Theorem 4.5.
Let $D$ be an adelic divisor with $D|_{T}=0$. The following two statements are equivalent.
(1) The adelic divisor $D$ is toric: the Green function $g_{D}$ is the pull back of th continuous conic function $SF_{D}:N_{\mathbb{R}}\to\mathbb{R}$ under the tropicalization map $\mathrm{trop}:T^{\mathrm{an}}\to N_{\mathbb{R}}$.
(2) There is a Cauchy sequence $(X_{i},D_{i})$ representing $D$ such that each $X_{i}$ is a complete toric variety and $D_{i}\in\mathrm{Div}(X_{i})$ is supported in the boundary $X_{i}\backslash T$.
Proof.
(2) $\Rightarrow$ (1): By Proposition 4.4 and the fact that the limit of toric adelic divisors are still toric.
(1) $\Rightarrow$ (2): We fix a projective model $X=\mathbb{P}^{n}$ of $T$, with boundary divisor $D_{0}=X\backslash T$, and consider the set of all functions $f:N_{\mathbb{R}}\to\mathbb{R}$ which are conical, continuous, piecewise linear, convex on each face, and has integral value on $N$. By [KKMS, Theorem 9], every such function gives a coherent sheaf of $T$-invariant fractional ideals contained in $i_{*}\mathscr{O}_{T}$, and then give a complete toric variety with a Cartier divisor on it whose supporting function is exactly $f$.
Take any norm on $N_{\mathbb{R}}$, and let $\Sigma$ be the unit sphere in $N_{\mathbb{R}}$. The restriction gives an isomorphism $C(N_{\mathbb{R}})_{\mathrm{eqv}}$ and $C(\Sigma)$. we will identify these two linear spaces.
Since the set of all such functions are closed under addition, taking minimum, and separating points in $\Sigma$ by the Boolean ring version of Stone-Weierstrass theorem, any continuous conical function can be approximated by these functions uniformly on all compact subset. Hence any toric adelic divisor can be approximated by model toric divisors. ∎
In fact, from the proof of Proposition 4.4, we see that $D$ is toric if and only if $D$ is $T$-invariant, meaning that for the multiplication and projection maps $m,p_{2}:T\times T\to T$, two adelic divisors $m^{*}D$ and $p_{2}^{*}D$ are equal with respect to the partial compactification $T\times T\to T\times X$. Since we don’t introduce such adelic divisors, we refer to [Son, §3.1].
4.4 Comparison with others
Comparison with toroidal $b$-divisors
Now we compare this theorem with the toroidal $b$-divisors in [BB, §4.3]. They introduced general toroidal embedding, but here we only consider the toric varieties to simplify the case. Then any allowable modification of a toric variety is given by birational toric morphism [KKMS, Page 24,90].
Let $X$ be a complete toric variety with open subset $U\cong T$. Let $\Sigma$ be the set of all proper birational toric morphism $X^{\prime}\to X$ with $X^{\prime}$ a smooth complete toric variety. Then in each $X^{\prime}$, the Cartier divisor and the Weil divisors are the same.
They define the toroidal Riemann–Zariski space as the inverse limit in the category of locally ringed spaces
$\mathfrak{X}_{U}:=\lim_{\begin{subarray}{c}\longleftarrow\\ X^{\prime}\in\Sigma\end{subarray}}X^{\prime}$
where $X^{\prime}\to X$ are all proper toroidal birational morphism w.r.t the subdivision of $N_{\mathbb{R}}$. Then they define the group of $\mathbb{R}$-Cartier toroidal $b$-divisors as
$\mathrm{CbDiv}(\mathfrak{X}_{U})_{\mathbb{R}}:=\lim_{\begin{subarray}{c}\longrightarrow\\ X^{\prime}\in\Sigma\end{subarray}}\mathrm{Div}(X,U)_{\mathbb{R}}$
with maps given by pull-back of Cartier divisors. It is endowed with inductive topology; and the toroidal Weil divisor as
$\mathrm{WbDiv}(\mathfrak{X}_{U})_{\mathbb{R}}:=\lim_{\begin{subarray}{c}\longleftarrow\\ X^{\prime}\in\Sigma\end{subarray}}\mathrm{Div}(X,U)_{\mathbb{R}}$
with maps given by push-forward of Weil divisors. It is endowed with the projective topology.
Proposition 4.6.
We have canonical injections
$\mathrm{CbDiv}(\mathfrak{X}_{U})\subset\widehat{\mathrm{Div}}(U)_{t}\subset\mathrm{WbDiv}(\mathfrak{X}_{U})$
Such that the corresponding supporting functions is given by
$\mathrm{PL}(N_{\mathbb{R}})\subset C(N_{\mathbb{R}})_{\mathrm{conic}}\subset\mathrm{Conic}(N_{\mathbb{R}}),$
where $\mathrm{PL}(N_{\mathbb{R}})$ is the space of piecewise linear conical functions on $N_{\mathbb{R}}$ whose linearity locus is defined over $\mathbb{Q}$, $C(N_{\mathbb{R}})_{\mathrm{conic}}$ is the space of continuous conical functions on $N_{\mathbb{R}}$, and $\mathrm{Conic}(N_{\mathbb{R}})$ is the space of arbitrary conical functions on $N_{\mathbb{Q}}$.
Be careful that their definition of $\varphi_{D}$ in [BB, Definition 4.2] differs with ours by a sign.
Proof.
The inclusion map $\mathrm{CbDiv}(\mathfrak{X}_{U})\subset\widehat{\mathrm{Div}}(U)_{t}$ is clear, and we only prove the second one $\widehat{\mathrm{Div}}(U)_{t}\subset\mathrm{WbDiv}(\mathfrak{X}_{U})$.
Suppose that the adelic divisor $D$ is represented by $(X_{i},D_{i})$, and all $X_{i}$ are complete normal toric varieties. We may assume that $D_{0}=X_{0}\backslash U$ is the boundary divisor. We also assume that for $j>i$, there exists a projection map $\pi_{ji}:X_{j}\to X_{i}$. Then there is a sequence $\epsilon_{i}\in\mathbb{Q}^{+}$, converging to 00, such that
$-\epsilon_{i}\pi^{*}_{j0}D_{0}\leqslant\pi^{*}_{ji}D_{i}-D_{j}\leqslant\epsilon_{i}\pi_{j0}^{*}D_{0}.$
Now let $Y\in\Sigma$. For each $i$, let $D^{\prime}_{i}$ be defined as follows. Take $p_{1}:Y\times_{X}X_{i}\to Y$ and $p_{2}:Y\times_{X}X_{i}\to X_{i}$. Then let $D^{\prime}_{Y,i}:=p_{1,*}p_{2}^{*}D_{i}$. It’s easy to check that $D^{\prime}_{Y,i}$ converges to some $E_{Y}$ in the finite dimensional space $\mathrm{Div}(Y,U)_{\mathbb{R}}$, and such that $\{E_{Y}\}$ form a Weil toroidal $b$-divisor. The rest part of supporting functions are easy. ∎
Comparison with special adelic divisors
We can also compare our definition with special adelic divisors in [Son, §5.]. We need to fix an smooth toric embedding $U\subset X$. Then each Cartier divisor on $X$ is also a Weil divisor. Let $D=X\backslash U$, and let $D=\sum D_{i}$ be the decomposition of $D$ into irreducible components. Since each $D_{i}$ is also Weil, there is a unique ray $L_{i}$ in $N_{\mathbb{R}}$ corresponding to $D_{i}$. Let $v_{i}\in N$ be a primitive generator of $D_{i}$.
We view $\operatorname{Spec}k((t))$ be the generic points in $\operatorname{Spec}k[[t]]$, then we can talk about order of adelic divisors on $\operatorname{Spec}k((t))$.
Definition 4.7.
An adelic divisor on $T$ is called special (with respect to $U\subset X$) if there is a continuous conic function $f:N_{\mathbb{R}}\to\mathbb{R}$ such that for all morphism $i:\operatorname{Spec}k((t))\to T$ with extension $i:\operatorname{Spec}k[[t]]\to X$, we have
$\mathrm{ord}_{0}i^{*}E=f\left(\sum_{j}\mathrm{ord}_{0}i^{*}D_{j}\cdot v_{j}\right).$
Note that we always have $\mathrm{ord}_{0}i^{*}D_{j}\geqslant 0$, and $\{j:\mathrm{ord}_{0}i^{*}D_{j}\neq 0\}=\{j:i([0])\in D_{j}\}$ has cardinality at most $n+1$, hence the sum is in fact on a simplicial cone.
Such a function $f$ is called the skeleton function in [Son].
Theorem 4.8.
An adelic divisor is toric if and only it is special with respect to a smooth toric compactification.
Proof.
Any smooth toric compactifications are essentially $\mathbb{G}m^{n}\subset\mathbb{A}^{n}$ locally, as it corresponding to a simplicial cones in $N_{\mathbb{R}}$ generated by a subset of $N$. Therefore, it suffices to consider the case that $X_{0}\subset\mathbb{A}^{n}$ and that $i:k[[t]]\to\mathbb{A}^{n}$.
Let $E$ ba a toric adelic divisor, and $i:\operatorname{Spec}k((t))\to T$ be a map which extends to $\operatorname{Spec}k[[t]]\to\mathbb{A}^{n}$. Take the valuation $|\cdot|_{s}$ on $k((t))$ such that $|t|=e^{-1}$. We also denote $s$ to be the unique point in $\mathcal{M}(k((t)),|\cdot|_{s})$, and $i^{\mathrm{an}}:\{s\}\to T^{\mathrm{an}}$ be its analytification. Then we have that $\mathrm{ord}_{0}i^{*}E=g_{E}(i(s))$.
If $E$ is toric, then the Green function $g_{E}(i(s))$ of $E$ at $i(s)$ is determined by its tropicalization
$\mathrm{ord}_{0}i^{*}E$ $=g(i(s))=f(\mathrm{trop}(i(s)))$
$=f\left(\sum_{j}\mathrm{ord}_{0}i^{*}D_{j}\cdot v_{j}\right)$
Therefore, $E$ is special.
Conversely, if $E$ is special, then we can find a toric adelic divisor with the same supporting function on $N_{\mathbb{R}}$, so it suffices to prove the case that if $E$ is special with supporting function zero, then $E$ is itself zero. This follows from Lemma 4.9 below. ∎
Lemma 4.9.
Let $U$ be any quasi-projective variety over $k$, and $D$ be an adelic divisor on $U$ with $D|_{U}=0$. Then $D=0$ if and only if for all smooth curves $C$ and maps $\lambda:C\to U$, the pull-back $\lambda^{*}D=0$.
Proof.
Suppose that the adelic divisor $D$ is represented by $(X_{i},D_{i})$, and all $X_{i}$ are normal. We may assume that $D_{0}=X_{0}\backslash U$ is the boundary divisor, and that for $j>i$, there is a morphism $\pi_{ji}:X_{j}\to X_{i}$. Then there is a sequence $\epsilon_{i}\in\mathbb{Q}^{+}$, converging to 00, such that
$-\epsilon_{i}\pi^{*}_{j0}D_{0}\leqslant\pi^{*}_{ji}D_{i}-D_{j}\leqslant\epsilon_{i}\pi_{j0}^{*}D_{0}.$
Now assume that $D\neq 0$. Then for $i$ large enough, the set
$\{E\in\mathrm{Div}(X_{i})_{\mathbb{Q}}:-\epsilon_{i}D_{0}\leqslant D_{i}-E\leqslant\epsilon_{i}D_{0}\}$
does not contain the zero divisor. In particular, there is a irreducible Weil divisor $F$ such that either $\mathrm{ord}_{F}D_{i}<-\epsilon_{i}\mathrm{ord}_{F}D_{0}$ or $\mathrm{ord}_{F}D_{i}>\epsilon_{i}\mathrm{ord}_{F}D_{0}$.
Now, choose a closed point $x\in X_{i}$ such that $x\in F$ but not in any other boundary irreducible divisors. Take a complete curve $C^{\prime}$ on $X_{i}$ passing through $s$ which is not contained in the boundary. Let $C$ be the normalization of $C^{\prime}\cap U$ and $\lambda:C\to X_{i}$ be the map. Then we can check that $\mathrm{ord}_{s}\lambda^{*}D\neq 0$. ∎
5 Global Monge-Ampère Measure
Let $k=\mathbb{Z}$ or a field, and let $U$ be a quasi-projective flat variety over $k$ of dimension $n+1$. Let $\widehat{\mathrm{Pic}}(U)_{\mathrm{int}}$ be the group of integrable adelic line bundles defined by Yuan–Zhang. They also defined a intersection theory of adelic line bundles in [YZ, §4], which is a symmetric multi-linear map on $\widehat{\mathrm{Pic}}(U)_{\mathrm{int}}^{n+1}$. It can be extend as follows.
Proposition 5.1.
The intersection pairing extends uniquely to a multi-linear map
$\widehat{\mathrm{Pic}}(U)\times\widehat{\mathrm{Pic}}(U)_{\mathrm{int}}^{n}\to\mathbb{R},$
such that if $\bar{L}_{1},\ldots,\bar{L}_{n}\in\widehat{\mathrm{Pic}}(U)_{\mathrm{int}}$ are fixed, then the map $\widehat{\mathrm{Div}}(U)\to\mathbb{R}$ given by $\bar{E}\mapsto(\bar{E}\cdot\bar{L}_{1}\cdots\bar{L}_{n})$ is continuous with respect to the boundary norm on $\widehat{\mathrm{Div}}(U)$. In other word, it is allowed to have at most one non-integrable adelic line bundle.
Proof.
Fix $\bar{L}_{1},\ldots,\bar{L}_{n}\in\widehat{\mathrm{Pic}}(U)_{\mathrm{snef}}$, and let $E\in\widehat{\mathrm{Div}}(U)_{\mathrm{mod}}$ be an effective model divisor. Then $\bar{E}$ is also integrable, and we have that
$(\bar{E}\cdot\bar{L}_{1}\cdots\bar{L}_{n})\geqslant 0.$ (1)
Now let $\bar{E}\in\widehat{\mathrm{Div}}(U)$ be an effective adelic divisor. Then $\bar{E}=\lim_{i\to\infty}\bar{E}_{i}$ for effective model adelic divisors $\bar{E}_{i}$. Fix a boundary divisor $\bar{D}_{0}$. Then there exist $\varepsilon_{i}\in\mathbb{Q}_{+}$ converging to 00, such that
$-\varepsilon_{i}\bar{D}_{0}\leqslant\bar{E}_{i}-\bar{E}_{j}\leqslant\varepsilon_{i}\bar{D}_{0}$
in terms of effectivities for all $j>i$.
Apply $(-\cdot\bar{L}_{1}\cdots\bar{L}_{n})$, by inequality (1), we have
$-\varepsilon_{i}(\bar{D}_{0}\cdot\bar{L}_{1}\cdots\bar{L}_{n})\leqslant(\bar{E}_{i}-\bar{E}_{j}\cdot\bar{L}_{1}\cdots\bar{L}_{n})\leqslant\varepsilon_{i}(\bar{D}_{0}\cdot\bar{L}_{1}\cdots\bar{L}_{n})$
Thus the sequence $(\bar{E}_{i}\cdot\bar{L}_{1}\cdots\bar{L}_{n})$ is a Cauchy sequence, and we define
$(\bar{E}\cdot\bar{L}_{1}\cdots\bar{L}_{n})=\lim_{i\to\infty}(\bar{E}_{i}\cdot\bar{L}_{1}\cdots\bar{L}_{n}).$
For general case, note that any adelic line bundles is the difference of two effective line bundles, we get our results. ∎
In particular, let $\widehat{\mathrm{Div}}(U)_{b}$ be the group of adelic divisors with zero integral part, then by Theorem 3.7, we have that $\widehat{\mathrm{Div}}(U)_{b}\cong C(U^{\mathrm{an}})_{\mathrm{eqv}}$, which is further non-canonically isomorphic to $C(\widetilde{U}^{b})$ for the compact Hausdorff space $\widetilde{U}^{b}$ defined in §2. We thus get a morphism
$\widehat{\mathrm{Pic}}(U)_{\mathrm{int}}^{n}\to(C(\widetilde{U}^{b}))^{\vee}\cong\mathcal{M}(\widetilde{U}^{b})$
where $\mathcal{M}(U^{b})$ is the Banach space of signed Radon measures on $U^{b}$, and the latter isomorphism is given by the Riesz representation theorem for measures. We call it the global Monge-Ampere measure associated to the integrable adelic line bundles $\bar{L}_{1},\ldots,\bar{L}_{n}$, denoted by $\mathrm{MA}(\bar{L}_{1},\ldots,\bar{L}_{n})$, and the intersection pairing is also a integration
$(\bar{E}\cdot\bar{L}_{1}\cdots\bar{L}_{n})=\int_{\widetilde{X}}h_{\bar{E}}(x)\ \mathrm{MA}(\bar{L}_{1},\ldots,\bar{L}_{n}),$
where $h_{\bar{E}}=g_{\bar{E}}/g_{\bar{D}_{0}}$ for the fixed boundary divisor $D_{0}$, and $\mathrm{MA}(\bar{L}_{1},\ldots,\bar{L}_{n})$ is also computed via $D_{0}$.
If all $\bar{L}_{i}$ are equal, the measure is denoted by $\mathrm{MA}(\bar{L})$.
To emphasize the independence of $\bar{D}_{0}$, we introduce the following notation.
$\int_{U^{b}}g(x)\mathrm{MA}(\bar{L}_{1},\ldots,\bar{L}_{n}):=\int_{\widetilde{U}^{b}}\frac{g(x)}{g_{\bar{D}_{0}}(x)}\mathrm{MA}(\bar{L}_{1},\ldots,\bar{L}_{n}),$
for any norm-equivariant measurable function on $U^{b}$, where the measure $\mathrm{MA}(\bar{L}_{1},\ldots,\bar{L}_{n})$ on $\widetilde{U}^{b}$ is computed via the same boundary divisor $\bar{D}_{0}$.
Question
Solve the equation $\mathrm{MA}(\bar{L})=\mu$ for a given semi-positive measure $\mu$ and a given line bundle $L\in\mathrm{Pic}(U)_{\mathrm{snef}}$.
We have the following naive conditions for existence and uniqueness of the Monge-Ampère equation.
Definition 5.2.
Let $\bar{E}\in\widehat{\mathrm{Div}}(U)_{b}$ be an integrable adelic divisor. Then $E$ is called numerically trivial on the boundary if
$(\mathscr{O}(\bar{D})\cdot\mathscr{O}(\bar{E})\cdot\bar{L}_{1}\cdots\bar{L}_{n-1})=0$
for all $\bar{D}\in\widehat{\mathrm{Div}}(U)_{b}$ and all integrable adelic line bundles $\bar{L}_{i}$.
We fix a model line bundle $(X,\bar{L}_{0})$ for $(U,L)$. Then any other line bundle $\bar{L}$ with $\bar{L}|_{U}\cong L$ can be written as a sum of $L_{0}$ with a boundary divisor.
Now consider the equation
$\mathrm{MA}(\bar{L})=\mu,$ (2)
which is equivalent to solve the equation $\mathrm{MA}(\bar{L}_{0}+f(x))=\mu$ for integrable norm-equivariant function $f(x)\in C(U^{b})_{\mathrm{eqv}}$.
Proposition 5.3.
(1) If the equation (2) has a solution, then $\bar{L}_{0}$ and $\mu$ satisfy the conditions
[…]
for all divisors $\bar{D}\in\widehat{\mathrm{Div}}(U)_{b}$ which are numerically trivial on the boundary.
(2) If $\bar{L}_{1}$ is a solution, then so is $\bar{L}_{1}+\mathscr{O}(\bar{E})$ for all adelic divisor $\bar{E}\in\widehat{\mathrm{Div}}(U)_{b}$ which are numerically trivial on the boundary.
The proofs are easy. We call them the naive conditions for the Monge-Ampère equation. It would be natural to ask whether these conditions are sufficient under some positivity conditions of $L$ and $\mu$.